### v1

# RoToR: Towards More Reliable Responses for Order-Invariant Inputs

Abstract
Mitigating positional bias of language models (LMs) for listwise inputs is a well-known and important problem (e.g., lost-in-the-middle). While zero-shot order-invariant LMs have been proposed to solve this issue, their success on practical listwise problems has been limited. In this work, as a first contribution, we identify and overcome two limitations to make zero-shot invariant LMs more practical: (1) training and inference distribution mismatch arising from modifying positional ID assignments to enforce invariance, and (2) failure to adapt to a mixture of order-invariant and sensitive inputs in practical listwise problems. To overcome, we propose (1) RoToR, a zero-shot invariant LM for genuinely order-invariant inputs with minimal modifications of positional IDs, and (2) Selective Routing, an adaptive framework that handles both order-invariant and order-sensitive inputs in listwise tasks. On the Lost in the middle (LitM), Knowledge Graph Question Answering (KGQA), and MMLU benchmarks, we show that RoToR with Selective Routing can effectively handle practical listwise input tasks in a zero-shot manner. [^1] https://github.com/soyoung97/RoToR
RoToR: Towards More Reliable Responses for Order-Invariant Inputs
Soyoung Yoon1thanks:   Work done during an internship at Channel Corporation.      Dongha Ahn1    Youngwon Lee1    Minkyu Jung2    HyungJoo Jang2    Seung-won Hwang1thanks:   Corresponding author. 1Seoul National University 2Channel Corporation {soyoung.yoon, seungwonh}@snu.ac.kr
1 Introduction
Figure 1: Self-attention alteration from order-invariant models. (a) PCW by elimination (b) PINE by re-assignment of position IDs based on query-based pairwise ordering. In contrast, (c) RoToR minimizes the distribution mismatch by global ordering with circular assignment.
Language conveys meaning in part through positional information, such as word placement and sentence structure. Given this nature, Language Models (LMs) that learn from human language are trained sensitive to positional information related to the ordering of segments. However, there are some listwise inputs that require neutrality to positional information. For example, for inputs such as sets, tables, databases, or multiple-choice questions, the ordering of the input segments—e.g., rows in a table or elements in an unordered set—require an order-agnostic understanding. We refer to such inputs as “order-invariant inputs,” which LMs reportedly struggle on. For example, in LLM-as-a judge scenarios, they exhibit a preference of up to 75% for the first answer in pairwise inputs Zheng et al. (2024b), and ranking between LMs can change up to 8 positions in different orderings of multiple choice questions on MMLU Alzahrani et al. (2024). Such results question the reliability of LM outputs on order-invariant inputs. Meanwhile, existing methods for enforcing invariance to LMs showed limited effectiveness in real-world tasks, which we hypothesize to arise from the following limitations.
First, training and inference distribution mismatch due to the positional ID re-assignment of zero-shot order-invariant LMs: Fig. 1 illustrates how self-attention is altered in there models. Unlike the original non-invariant model which always assigns position IDs in a causal, ascending manner, order-invariant models either eliminate inter-segment attention, such as PCW Ratner et al. (2023) in Fig. 1a, or re-assign position IDs as in PINE Wang et al. (2024) in Fig. 1b, re-ordering segments using an importance score, placing similar segments closer to the query. For each query segment, it computes segment-wise query-key attention (for each attention head in each decoder layer) and re-assigns position IDs of segments as keys. This query-dependent segment ordering leads to excessively frequent alterations on positional ID assignments. Frequent re-assignments can also confuse the model and risk collisions which violates the invariance property (e.g., multiple segments having the same importance score).
To overcome this, we propose a query-agnostic global sorting with circular arrangement for order-invariant positional ID assignment. Ours is named RoToR, inspired by the word rotary to express circular assignment, and also a palindrome, to reflect order invariance. Fig. 1c contrasts with PINE in Fig. 1c, where RoToR only needs a single global ordering (e.g., A->B->O->K->G) with no extra attention computation. The ordering of segments on suffix tokens remain in a fixed order, since it does not rely on their similarity to the query. Finally, we propose three different global sorting algorithms for RoToR, and demonstrate that they consistently outperform previous order-invariant models.
Second, in practical listwise inputs, order-invariant task may include order-sensitive inputs which requires order-specific understanding. For example, the (d) None of the above option in MMLU cannot be reordered. For such a “mixed” nature requires to handle each of them adaptively by a simple Selective Routing method. Selective Routing adapts to input by routing between two models, invariant and non-invariant (original), based on their confidence scores of predictions. Experiments on the MMLU benchmark show that Selective Routing effectively handles datasets with order-invariant and sensitive inputs, and achieve better order robustness while maintaining the original performance.
In summary, our contributions are as follows: 1. Clarifying key challenges to robust understanding of listwise inputs. We pinpoint the distribution mismatch and positional ID assignment complexities that hinder zero-shot order-invariance in LMs, and the need to adaptively handle order-invariant and order-sensitive inputs. 2. A stable, order-invariant solution (RoToR): We propose a query-agnostic global ordering with minimal positional ID modifications, resulting in stable and efficient order-invariance. 3. Adaptive handling of listwise inputs (Selective Routing): We introduce a simple routing method that switches between the original and invariant LMs based on confidence. On MMLU, we show that Selective Routing can adaptively deal with both types of input, leading to better stability. To this end, we aim to develop a model that excels at processing a wide range of listwise inputs reliably and efficiently.
2 Related Works
2.1 Positional bias of LLMs
Problem statement. Recent works on (zero-shot) retrieval augmented generation (RAG) with LLMs have found that the models exhibit unwanted bias on the ordering of the retrieved documents Chhabra et al. (2024). Widely known as the lost-in-the-middle problem Liu et al. (2024), many prior studies Chen et al. (2024); Gupta et al. (2024); Pezeshkpour and Hruschka (2023); Zhao et al. (2023); Zhou et al. (2024); Wei et al. (2024); Alzahrani et al. (2024); Zheng et al. (2024a) also investigate the importance of positional bias, extending the domain to structured knowledge grounding (SKG) tasks Zhao et al. (2023); Zhou et al. (2024) and multiple-choice questions Gupta et al. (2024) where changing the ordering of rows, schemas, or choices greatly degrades performance.
Considerations for decoder-only LMs. While successful approaches are presented to mitigate this issue for encoder-only Yang et al. (2022) and encoder-decoder Yen et al. (2024); Cai et al. (2023) models, leaving decoder-only models, which account for most of the current LLMs, for more consideration. In contrast to transformer encoders that use bidirectional attention which is invariant by nature Lee et al. (2019), transformer decoders use causal attention in order to learn causal relation signals, which is not invariant by nature Haviv et al. (2022a). Therefore, positional bias for decoder-only models is known to stem from both positional encoding and causal attention mask Yu et al. (2024); Wang et al. (2024).
2.2 Zero-shot order-invariance for LLMs
Long context modeling. Zero-shot approaches for mitigating positional bias in LLMs were first raised in long-context tasks, with a goal to correctly handle relevant information located in the middle of lengthy inputs222github.com/gkamradt/LLMTest_NeedleInAHaystack. Nonetheless, these works focus primarily on understanding long texts without losing precision Li et al. (2023); Zhang et al. (2024a); An et al. (2023); Bai et al. (2024), whereas positional bias is a more general problem that can occur even on multiple-choices questions with relatively short contexts Alzahrani et al. (2024). Technically, this line of works modify the attention mechanism by altering the positional encoding to adapt an LLM to longer contexts Peng et al. (2023); Hsieh et al. (2024); Peysakhovich and Lerer (2023); Chen et al. (2023); Junqing et al. (2023); Xu et al. (2023); Yu et al. (2024); Zhang et al. (2024b). But since they do not modify the causal mask which also contributes to positional bias, order-invariance is not guaranteed in general Haviv et al. (2022b).
(Zero-shot) order-invariance. Recent line of works focused on achieving order-invariance by mechanistically altering both positional encoding and causal masking. While several works require training Junqing et al. (2023); Zhu et al. (2023), we focus on zero-shot approaches for practicality, namely PCW, Set-Based Prompting Ratner et al. (2023); McIlroy-Young et al. (2024), and PINE Wang et al. (2024), which we explain in detail at Sec. 3.1. Another line of works based on self-consistency try to mitigate positional bias simply by running inference multiple times with different orderings of contexts Zheng et al. (2024a). However, in principle, this requires evaluating $n!$ forward passes in total, enforcing Monte Carlo approximations (Tang et al., 2024). More recent work optimizes the number or passes (Lee et al., 2025b) with similar comprehensiveness (Hwang and Chang, 2007), or replaces with contrastive training objective Lee et al. (2025a). In contrast, our method guarantee invariance with a single forward pass, without requiring any approximations.
3 Methodology
Figure 2: For example input ‘‘Given Apple Banana Orange which one is most related to the color red? A: It is ...’’, demonstration of how PINE and RoToR (ours) modify the attention mask and positional IDs for each query-key combination, resulting in segment-wise order invariance. Assume each block represents a single token, and blocks are colored in positional ID assignment. White (no-block) area indicate masked attention.
3.1 Baseline: Order-invariant causal LMs
In this section, we briefly overview the existing work on turning decoder-only LLM architectures into order-invariant models by adjusting attention mechanism, and review their limitations.
Isolated parallel processing
Prior works like PCW (Ratner et al., 2023) and Set-Based Prompting McIlroy-Young et al. (2024) have modified the attention mask and positional ID assignments of the language model to isolate the processing of each segment and apply same positional embeddings are applied across segments, and thus achieve order invariance: However, this design completely prevents one segment from attending to the others, and aggregating the information from different segments is solely handled at suffix and generated tokens, significantly hindering the LM’s cross-segment contextualized understanding of the text. Yang et al. (2023) have argued that this essentially degenerates to mere ensemble of conditioning on each context separately. Such information bottleneck and train-test time discrepancy limits the applicability, more severely as the number of segments is increased.
Bidirectional processing with Q-K similarity
A more recent work, PINE (Wang et al., 2024) has addressed these issues through a bidirectional attention mechanism, by letting each segment attend to all other segments. However, to allow this within decoder-only models with causal attention and still achieve order invariance, PINE modifies the multi-head self-attention procedure to create an ‘illusion’: it treats each query segment as if it were the final segment in the input list, so that it can attend to all other segments as keys. For each query segment, the ordering among the key segments is determined by their attention scores (without positional embeddings, $\mathrm{Attn}_{\mathrm{NoPoS}}$), ensuring that segments with stronger relevance to the query appear closer. Meanwhile, the tokens within a query segment still follow causal ordering. Fig. 2 illustrates this with the input [T1‘‘Given’’, S1[‘‘Apple’’], S2[‘‘Ban’’, ‘‘ana’’], S3[‘‘Orange’’], T6‘‘which one’’, .. T10‘‘red?’’]. Prefix tokens ‘‘Given’’ and suffix tokens ‘‘which one .. red?’’ remain in their original positions and follow normal causal attention, equally attending to all segments. In contrast, the segments in the middle (S1 - S3) require the order-invariant mechanism. PINE dynamically re-assigns positional IDs depending on a token’s role as a query or a key, instead of using a single fixed positional ID. For Case 1: When “ana” (T4) in segment S2 acts as a query, S2 is assigned the largest position IDs (5) so it can attend to all previous segments. Within S2, the tokens “ban” (4) and “ana” (5) still maintain their local causal order (i.e., smaller to larger positions). Case 2: When “ana” serves as a key, the placement of S2 depends on its $\mathrm{Attn}_{\mathrm{NoPoS}}$ scores with the query segment. If “banana” (S2) is more relevant to the query segment “apple” (S1) than “orange” (S3), (i.e., $\mathbf{Attn_{NoPos}}$(S1, S2) > $\mathbf{Attn_{NoPos}}$ (S1, S3)) S2 is placed nearer to S1, with “ban” and “ana” preserving their internal sequence. For prefix, suffix, and generated tokens, they do not participate in order invariance, so they are always placed at their standard causal positions. However, the same dynamic reordering applies to segment tokens (T2–T5) when computing attention scores of suffix (T6–T10) and the generated token (T11) as query.
Figure 3: Comparing the ordering of 5 segments (S1 - S5) of PINE (Wang et al., 2024; left) and RoToR (Ours; right). PINE sorts segments using aggregated attention scores. In order to be fully ordering-invariant, segment sorting is changed per token in suffix level, causing confusion. In contrast, we define one global sorting of segments and conduct circular assignment between segments. With this, we simply use the global sorting for position id assignment on suffix tokens, without harming invariance.
3.2 RoToR: minimal OOD from positional ID assignments
While PINE achieves order-invariance by contextualization across segments, its query-specific ordering scheme introduces (1) significant train-test behavior discrepancy as well as (2) unnecessary complexity and numerical instability, which limits its scalability. During decoding with PINE, position IDs are assigned differently for every query token (each token in the suffix), decoder layer, and attention head, as the query-key attention score $\mathrm{Attn}_{\mathrm{NoPos}}$ determines the position IDs. This complexity introduces excessively frequent alterations on position IDs: As the base LM is trained with fixed positional IDs and causal masks, this causes hidden activations higher risk of out-of-distribution (OOD) for it to process properly. Moreover, ordering segments based on attention is computationally expensive and introduces numerical instability. As computing the attention value of one query segment requires computing the KV attention over every other number of segments, PINE invokes $\mathcal{O}(n^{2})$ cost overhead for each segment for input length $n$, which is further multiplied by the number of all combinations of layers, heads, and the number of suffix and generated tokens. Also, in practice, calculating attention without RoPE results in a very narrow range of values. bfloat16 numeric type lacks precision to distinguish these values, leading to non-determinism originating from several tied values. The outcome may depend on the initial ordering; to address these problems stemming from query-dependent ordering, we instead propose RoToR (Fig. 3), which uses one global ordering that is not a function of the initial ordering of segments (e.g., canonical ordering by lexical sorting) and assign IDs for tokens in different segments based on circular arrangement.
Global ordering
Instead of re-computing the relative order of segments for each query, we reuse a globally shared single ordering, avoiding costly recomputation of numerically unstable attention scores. Moreover, this further reduces the gap between the LLM’s pretrained behavior and test-time behavior, as consistent position IDs are assigned across layers/heads/across suffix tokens. Global ordering allows to preserve the relative placement of segments, further closing the gap induced from introducing order invariance to causal LMs. For example, in Figure 3, due to the global ordering, segments $S_{5}$ and $S_{2}$ are always placed in adjacent positions with RoToR (right side), while it is not satisfied and constantly changed with PINE (left side). We consider three separate global sorting algorithm to be used in RoToR: (1) simple lexicographical sorting which can be obtained with minimal overhead based on tokenized sequence of segments, (2) using a pointwise reranker333castorini/monot5-base-msmarco-10k to score relevancy of each row with respect to the question, or (3) simple frequency-based sorting which normalizes token ids based on the inverse frequency of each token (Details at Fig. 6). Empirically, we find that using simple lexicographical sorting is sufficient for bringing improvements over PINE.
Circular arrangement
To mimic bidirectionality with causal LMs, each segment should be assigned position IDs so that they appear to themselves as being placed at the end of the sequence of segments. To achieve this with a shared global ordering, we employ circular arrangement, each segment taking turns to be placed at the end while their relative ordering is preserved. Given the global ordering, we can construct a single directed graph by combining the front and last parts. Then, we assign orderings for each segment as query by following the path from the graph, starting from the query segment, which is illustrated in Fig. 3. For all suffix and generated tokens, segments are arranged according to the initial front and last part of the global ordering. Compared to PINE where we have to assign different orderings of segments for each suffix and generated tokens, RoToR assign the same positional ID, acting merely the same as the original token. This also accounts for reducing the distributional gap between the original model.
Computational overhead
By adopting global sorting, we can avoid extra attention scores computation and thus improve efficiency. While the cost with PINE to obtain hidden states is $\mathcal{O}(n^{2}d+nk\log{k})$ for input length $n$, hidden dimension $d$, and $k$ segments (Wang et al., 2024), our method with lexicographical sorting achieves $\mathcal{O}(nk\log{k})$ cost, being more efficient and faster (lightweight) than PINE. We empirically validate that our method is much faster than PINE as $k$ increases, at Tab. 4.
3.3 Selective Routing for handling order-sensitive inputs
Figure 4: Illustration of Selective Routing (Sec. A.5).
Since many practical benchmarks such as MMLU involves semi-invariant inputs, we propose a routing mechanism that uses the order-invariant model in conjunction with the standard causal model for further applicability. Our design is partly based on the finding from Wei et al. (2024) that there is correlation between task difficulty (which is in turn correlated with confidence values) and the model’s sensitivity to ordering. Selective Routing, illustrated in Fig. 4, combines confidence, the model output probability for the generated answer, from two different model versions—the original model and the order-invariant model—on the same input and choose a more confident answer. Both models first produce a maximum probability over possible answer tokens (e.g., A, B, C, D for MMLU) and a corresponding answer choice. We then compare the original model’s maximum probability, plus a bias term $\alpha$, to the invariant model’s maximum probability. If the original model’s adjusted score is higher, we take its answer; otherwise, the invariant model’s answer is chosen. $\alpha$ is a hyperparameter that controls how strongly the original model is favored, which was selected as 0.2 according to hyperparameter search on the validation subset (Appendix Sec. A.5).
4 Experiment setup
Total ndoc (segments) 10 20 30
Gold idx at: 0 4 9 avg. 0 4 9 14 19 avg. 0 4 9 14 19 24 29 avg.
Indexing bias present
Original 52.8 51.1 46.7 50.2 55.0 50.7 50.5 50.2 47.5 50.7 56.1 50.8 51.5 51.1 50.8 50.5 48.9 51.4
PCW 12.0 11.9 12.1 12.0 03.4 03.7 03.8 03.9 03.6 05.1 02.1 02.1 02.0 01.9 02.0 02.2 02.0 02.0
Set-Based Prompting 40.8 40.7 40.8 40.8 25.6 25.8 25.7 25.5 25.3 25.6 15.8 15.9 16.1 16.0 16.1 15.7 15.8 15.9
PINE 59.2 56.8 57.7 57.9 56.2 55.1 54.5 54.2 55.3 55.5 56.1 53.6 53.3 53.9 53.2 53.7 54.5 54.0
RoToR-lexical 63.6 60.5 59.1 61.1 52.6 58.6 57.8 56.8 58.6 57.6 64.6 58.9 56.9 56.2 57.1 56.2 57.1 58.1
RoToR-reversed lexical 61.5 60.8 60.6 61.0 60.8 58.6 59.7 60.5 59.3 60.0 61.1 57.5 58.2 58.6 58.3 59.3 59.2 58.9
RoToR-reranking 61.4 61.4 61.7 61.5 62.3 59.2 59.1 59.4 59.8 60.2 62.2 58.6 58.4 58.6 58.3 58.5 59.8 59.2
RoToR-freq 62.8 61.1 59.5 61.1 62.9 58.8 56.7 57.4 58.0 59.1 61.7 58.2 56.9 56.1 56.4 55.4 56.8 57.4
Indexing bias removed
Original 54.7 53.0 50.2 52.6 54.8 52.6 52.8 52.4 51.0 52.7 55.6 51.5 52.4 52.8 52.1 52.3 53.0 52.8
PCW 12.4 11.9 12.2 12.2 03.7 04.0 04.0 04.0 03.9 03.9 02.3 01.8 02.0 02.0 02.1 02.0 02.0 02.0
Set-Based Prompting 42.5 42.5 42.5 42.5 26.3 26.3 26.3 26.3 26.3 26.3 14.1 14.1 14.1 14.1 14.1 14.1 14.1 14.1
PINE 58.6 58.8 59.0 58.8 56.2 55.7 55.5 55.7 55.5 55.7 54.2 54.8 54.3 53.7 54.8 54.2 54.0 54.3
RoToR-lexical 61.4 61.6 61.6 61.5 61.4 59.8 59.6 59.6 59.8 60.0 59.2 59.5 59.4 59.1 59.0 59.3 59.1 59.2
RoToR-reversed lexical 61.6 61.8 61.8 61.8 58.9 59.3 58.8 58.6 58.7 58.8 57.9 58.2 57.9 57.4 57.9 57.6 57.5 57.8
RoToR-reranking 61.2 61.4 61.2 61.3 60.9 61.0 61.2 61.2 61.2 61.1 60.9 60.7 60.7 60.7 60.8 60.8 60.7 60.8
RoToR-freq 61.0 61.1 61.1 61.1 60.4 60.3 58.6 60.2 60.0 59.9 59.3 60.4 59.7 59.5 59.5 59.6 59.2 59.6
[…]
WebQSP (Count = 1466) Mintaka (Count = 4000)
N=30 N=50 N=30 N=50
Acc EM F1 Avg. Acc EM F1 Avg. Acc EM F1 Avg. Acc EM F1 Avg.
Initial (no shuffling of segments)
Orig. 71.3 43.2 58.3 57.6 71.6 43.0 58.1 57.5 50.2 44.0 51.9 48.7 50.0 44.0 51.7 48.6
PINE 72.0 44.1 59.0 58.4 72.4 44.4 58.5 58.4 51.5 45.0 52.6 49.7 51.6 45.1 52.6 49.8
RoToR 73.7 43.4 59.2 58.8 73.9 43.5 59.0 58.8 53.1 46.5 54.1 51.2 52.9 46.0 53.6 50.8
RoToR (MonoT5) 72.9 43.6 59.2 58.6 73.9 43.9 59.1 59.0 51.6 45.0 52.5 49.7 52.4 45.4 52.8 50.2
RoToR (Freq) 73.5 43.5 59.3 58.8 74.1 43.9 59.3 59.1 52.6 46.1 53.7 50.8 53.1 46.4 53.7 51.0
After shuffling segments, averaged over 3 seeds
Orig. 70.9 42.4 57.8 57.0 $\pm$ 0.5 70.7 41.6 57.2 56.5 $\pm$ 0.2 49.5 43.3 51.0 47.9 $\pm$ 0.0 49.7 43.5 51.0 48.1 $\pm$ 0.4
PINE 71.9 43.9 59.1 58.1 $\pm$ 0.0 72.5 44.6 58.5 58.5 $\pm$ 0.3 51.8 45.2 52.8 50.0 $\pm$ 0.1 51.8 45.3 52.7 49.9 $\pm$ 0.2
RoToR 73.6 43.2 59.1 58.7 $\pm$ 0.1 74.0 43.3 58.8 58.7 $\pm$ 0.1 52.8 46.2 53.8 50.9 $\pm$ 0.0 52.7 45.9 53.5 50.7 $\pm$ 0.0
RoToR (MonoT5) 72.6 43.4 59.2 58.3 $\pm$ 0.1 73.9 43.7 58.9 58.8 $\pm$ 0.1 51.6 45.0 52.6 49.7 $\pm$ 0.1 52.2 45.2 52.8 50.1 $\pm$ 0.2
RoToR (Freq) 72.9 43.2 59.0 58.4 $\pm$ 0.0 73.7 43.8 59.1 58.9 $\pm$ 0.1 52.5 45.9 53.5 50.6 $\pm$ 0.1 53.1 46.4 53.7 51.1 $\pm$ 0.0
Table 2: Knowledge Graph Question Answering (KGQA) results, with standard deviation of the average scores with $\pm$. N refers to number of top-k segments per query.
Initial
ordering
Reversed
ordering
Avg. of all possible
(4!-1) re-orderings
Baseline: Single use of order-invariant models
Orig. 68.3 64.8 65.5 $\pm$ 1.0
PCW 57.0 55.1 56.1 $\pm$ 1.1
Set-Based Prompting 31.1 33.0 31.6 $\pm$ 0.8
PINE 64.8 63.3 63.6 $\pm$ 0.7
RoToR 63.2 62.6 62.8 $\pm$ 0.7
RoToR (MonoT5) 64.2 62.9 63.5 $\pm$ 0.5
RoToR (Freq) 64.3 63.5 63.6 $\pm$ 0.6
With Selective Routing
Orig. 68.3 64.8 65.5 $\pm$ 1.0
Orig + PINE 68.5 65.1 65.7 $\pm$ 0.9
Orig + RoToR 68.5 65.0 65.7 $\pm$ 0.9
Orig + RoToR (MonoT5) 68.4 65.2 65.8 $\pm$ 0.9
Orig + RoToR (Freq) 68.5 65.3 65.8 $\pm$ 0.8
Orig + RoToR (Optimal) 75.0 71.9 72.7 $\pm$ 1.0
Table 3: Improving applicability to general listwise tasks (MMLU, N=4) with Selective Routing, which includes both order-invariant and order-sensitive examples. RoToR with Selective Routing shows improved performance and stability across re-orderings of input, and its high upper-bound (Optimal) implies the potential for further improvements.
4.1 Baselines
Original causal LM with no modifications (Orig.) were compared, which processes text sequentially. Also, we compare RoToR against previous zero-shot order-invariant LMs discussed in Sec. 3.1, namely PCW Ratner et al. (2023), PINE Wang et al. (2024), and Set-Based Prompting McIlroy-Young et al. (2024). We use the LLaMA 3.1 AI (2024) 8B-Instruct444meta-llama/Meta-Llama-3.1-8B-Instruct as our backbone model for all of our experiments [^5] Results on 70B model is available in Appendix, Tab. 5.. As our method doesn’t need training, a single A6000 GPU was sufficient to run all of the experiments with our primary backbone model.
4.2 Benchmarks with listwise inputs
We experiment with three benchmarks involving real-world listwise input data. Examples of exact inputs and outputs are provided in Appendix A.7. All reported scores are rounded to the nearest tenth.
Knowledge Graph Question Answering (KGQA)
In KGQA tasks, the model takes facts over knowledge graphs represented as (subject, relation, object), and answers the given question based on the given facts. We basically follow the KGQA dataset preprocessing and evaluation setup from  Baek et al. (2023), which uses WebQSP Yih et al. (2016) modified from  Berant et al. (2013) and Mintaka Sen et al. (2022) with Wikidata for knowledge source, and use the Exact Match (EM), Accuracy (Acc), and F1 score metric for evaluation. We also use MPNet Song et al. (2020) as a dense retriever to retrieve top-k facts over each question, and experiment with segment size of 30 and 50. Replication details and example dataset format are at Appendix Sec. A.3, Fig. 9 and Fig. 10, respectively. Along with measuring the performance of the initial input ordering, we report performance after we shuffle the order of segments with 3 different seeds to see shuffle robustness.
Lost in the middle (LitM) We use the Lost in the Middle (LitM) benchmark Liu et al. (2024), which draws from 2655 queries in the Natural Questions (NQ) dataset. It provides sets of (10, 20, 30) passages, placing the gold passage at predetermined positions (e.g., 0, 4, 9) and filling the remaining slots with irrelevant passages, testing the model’s robustness to passage ordering. Following  Liu et al. (2024), we measure answer quality with the best_subspan_em metric. Experiments on LitM found that eliminating the effect of index bias is another important detail for invariant models to be effective: (Appendix Sec. B) thus, we also experiment with a variant of the benchmark with index bias eliminated (prompt at Appendix Fig. 8) in order to further investigate the impact with and without index bias.
MMLU The Massive Multitask Language Understanding (MMLU) benchmark Hendrycks et al. (2021) consists of 57 diverse sub-tasks with a total of 14,015 queries to measure general performance of LMs about the knowledge of the world. Despite its popularity, many works report that performance fluctuates heavily depending on the order of choices Gupta et al. (2024); Pezeshkpour and Hruschka (2023); Wei et al. (2024); Alzahrani et al. (2024); Zheng et al. (2024a) and is widely investigated to measure the positional bias of the model. On experiments on MMLU, we notice that a lot of proportions consist of ordering-sensitive inputs, which showed the effectiveness of adaptively applying Selective Routing. In addition to measuring the performance with the original ordering, we report the average performance for all possible (4!-1) re-orderings.
4.3 Further comparison with PINE
We additionally conduct experiments on a subset of benchmarks (LitM and MMLU) on the runtime latency, perplexity, and collision rate of PINE and RoToR, to validate our claims on Sec. 3.2.
5 Results & Analysis
Overview
We show results for KGQA in Tab. 2, and results for MMLU in Tab. 3. Results for LitM are in Tab. 1, with a visualization in Appendix Fig. 5. We use lexical sorting for RoToR unless stated otherwise.
Impact of removing index bias on LitM
On LitM, results in Tab. 1 show that, while invariant LMs show stable performances regardless of the gold index when index bias is removed (as described in Sec. 4.2) (Appendix Fig. 5(b)), fluctuations are observed if index bias is present (Appendix Fig. 5(a)). This implies that index bias serves as an implicit source of additional positional bias, and invariant LMs benefit from removing it.
Effectiveness of RoToR
We observe that shuffling input segments leads to non-trivial performance degradations in the original model, which exhibits a statistically significant performance drop on our experimented dataset (one-tailed t-test, p < 0.05). In contrast, our proposed RoToR model does not show a statistically significant difference in performance before and after shuffling, indicating that it is more robust against such perturbations (Appendix C). On LitM (Tab. 1), we notice PCW and Set-Based Prompting has impractical performance, with PINE degrading heavily (-4.5%) as number of documents ($k$) increases, while RoToR is less affected. On KGQA (Tab. 2), we show RoToR outperform PINE with lower standard deviation across shuffled segments. The effectiveness increases as N increases (30 -> 50). Practical scalability with increasing $k$ is critical, but we find that previous order-invariant LMs struggle handling larger $k$ (on KGQA and LitM). In contrast, RoToR shows better performance with improved efficiency and robustness.
Improvements from PINE
Runtime latency (s)
Perple-
xity ($\downarrow$)
Collision
Rate($\downarrow$)
MMLU LitM LitM
data count 14015 2655 2655
# segments 4 10 30 20 30
PINE 7,371 18,551 41,664 6.91 42.3%
RoToR 6,608 14,264 23,569 6.65 0 (None)
Reduced % 10.4% 23.1% 43.4%
Table 4: Compared to PINE, RoToR with lexical sort is faster, has lower perplexity, and have zero collision rate between segments. We use a subset of LitM with index bias removed and gold index = 0.
Experiments against comparing RoToR with PINE (Tab. 4) analyze the following: Runtime: Actual inference times (Appendix Sec. A.6) and find that RoToR outperforms PINE substantially, with efficiency gains increasing alongside $n$. For instance, on LitM (30 docs), RoToR achieves a 43% reduction in total runtime. Perplexity: Lower generation perplexity indicates input representations are closer to in-distribution. On the same LitM dataset, RoToR’s reduced perplexity implies its positional ID assignment effectively mitigates out-of-distribution effects. Collision Rate: PINE’s similarity-based ordering often collides: on average, only 17.3 of 30 similarity values are unique, causing 42% of the segments to be indistinguishable and thus breaking invariance. In contrast, RoToR with lexical sorting only collides if the segment texts are identical. On LitM, this yields zero collisions, preserving full invariance.
Selective Routing
Selective Routing on MMLU (Tab. 3) is a representative of a task that involves not only order-invariant, but also order-sensitive (e.g., "None of the above"), inputs. Therefore, single use of order-invariant models does not always outperform the original model, limiting applicability of order-invariant models to practical listwise tasks. For example, we observe significant performance drops for Set-based Prompting in MMLU, falling short of half the performance of the original model on initial ordering. Using RoToR with Selective Routing to handle order-sensitive inputs can be a solution to this case, which outperforms the original model in all possible orderings of candidate choices. Selective Routing improves the generalizability and extends the applicability on practical listwise tasks by adaptively handling order-sensitive inputs.
We additionally report the the RoToR + Selective Routing (Optimal) performance on Tab. 3, which improves by a significant 7% over using a single model. It was evaluated using a relaxed accuracy metric based on the union of predictions from the original and the invariant (RoToR-lexical) model. This highlights the potential of Selective Routing for further accuracy gains through optimizing design choices on routing methods, which we plan to explore in future work.
Impact of global ordering algorithm
While most of our experiments focus on the simplest lexical sorting method, RoToR supports any global sorting approach. To demonstrate this flexibility, we report experiments with various global sorting strategies, including reversed lexical sorting, MonoT5-based reranking, and token frequency-based sorting. Lexical sort is presented as a baseline (lower bound) - a simple algorithm ensuring global sorting. Our experiments on Tab. 1 show that any type of global sorting, with the use of circular assignment is superior than PINE, which relies on pairwise attention arrangements.
6 Conclusion
Our work addresses order-invariance in listwise inputs by identifying core issues in distribution mismatch and adaptive handling of mixed inputs. Our proposed RoToR provides a stable zero-shot order-invariant solution that reduces the complexity of positional ID modification, while MoV adaptively routes between invariant and sensitive LMs to handle real-world scenarios. Together, these methods demonstrate improved performance and reliability on KGQA, LitM, and MMLU benchmarks.
7 Limitations
Our method can utilize any kind of deterministic sorting algorithm, but we have only experimented with limited global sorting algorithms due to time and resource constraints. We plan to investigate potentially better sorting algorithms in the future. Also, current ordering-invariant models are limited to inputs given as prefix + parallel + suffix. It would be beneficial to support more complex structures, such as ability to process multiple order-invariant contexts interleaved with serial text.
Appendix A Appendix
A.1 Exact values of the lost-in-the middle benchmark
(a) Results with index bias (indexed by numbers). (Example input at Appendix Fig. 7)
(b) Results without index bias (indexed by title) (Example input at Appendix Fig. 8)
Figure 5: Results on the Lost-in-the-middle benchmark. Visualization of the best_subspan_em results at Appendix Tab. 1. RoToR (dark red, red, yellow) generally performs the best regardless of the position of the gold index, with less fluctuations when we remove index bias. Ours is RoToR with lexical sort, and ours-reversed is the one with the reversed lexical ordering. For brevity, only the performance of RoToR with reranking sort (MonoT5) is annotated as numbers, and the performance of PCW and Set-Based Prompting are reported only at the Table (Appendix Tab. 1) due to its low performance.
ndoc = 10 20
gold idx = 0 4 9 avg. 0
Orig. 66.2% 65.7% 65.7% 65.9% 65.2%
PINE 67.9% 67.8% 67.5% 67.7% 65.9%
RoToR-lex. 69.6% 69.5% 69.3% 69.5% 67.6%
Table 5: Reporting the performance with Llama-3.1-70B-Instruct on a subset of lost-in-the middle benchmark (without index bias) performance. We see that the gains are consistant with the models we mainly investigated (Llama-3.1-8B-Instruct)
Tab. 1 in the main paper reports the full number to the lost in the middle benchmark, and Fig. 5 visualizes them in graph format. Table  5 shows results in different model (Llama-3.1-70B-Instruct).
A.2 Illustration of the global sorting method
Figure 6: Illustration of ordering 7 rows by 2 different global sort options, (1) lexical sort based on token ids, or (2) reranking sort based on a reranker model (MonoT5-base in this case). (3) frequency based sorting applies lexical sorting, but the definitive token ids are mapped into the inverse frequency.
We show the two different global sorting algorithms presented in our paper at Fig. 6.
A.3 Details about preprocessing and evaluation of datasets
A.3.1 General.
All inferences were done with a single NVIDIA RTX A6000 48GB GPU. For reproducibility, we fix the seed and use greedy decoding without any sampling for all experiments. For all datasets we tested, we separate the input to 3 parts: prefix, parallel contexts, and suffix and feed them accordingly to the positional-invariant model. For the case of the original model, we simply join the prefix, context, and suffix text to make one sequential text. For testing with the PCW model, we concat each prefix to the parallel context due to the architectural limitations of the PCW model, which doesn’t have the prefix part (which is processed casually before parallel contexts). For example, on testing the PCW model on MMLU, we append the question and each answer options, which generally result in longer input sequences. For PCW, we use the pcw_generate function, and we additionally utilized the RestrictiveTokensLogitsProcessor provided at the official PCW repository for MMLU classification, to have a similar setup with the log_likelihood option used for other models.
A.3.2 Knowledge Graph Question Answering
For evaluation with Mintaka and WebQSP, we follow the same setup as  Baek et al. (2023). Given the gold answer and model generated answer, the EM score counts if both are exactly the same; Accuracy measures if the generated answer includes the gold answer, and F1 score measures the precision and recall among overlapping words. Since we are testing on a non-trained zero-shot version of the model, we enforce the model to output in json format to make it easier to parse. For the row shuffling setup, we fix the seed to 0, 1, 2 on shuffling rows and report the average scores.
A.3.3 Lost in the Middle
Specifically, we use the dataset provided in the official repository666github.com/nelson-liu/lost-in-the-middle, use the same prompt as the llama 2 chat model with only the instruction tokens adjusted to llama 3 (removed [Inst] and changed to <|begin_of_text|> and etc.,), and evaluate using the best_supspan_em metric.
A.3.4 MMLU
We follow the publicly acknowledged lm-evaluation harness Gao et al. (2024) prompt design by eluther.ai. We measure accuracy between the gold answer and the token with the highest likelihood (probability) among possible answer tokens [‘ A’, ‘ B’, ‘ C’, ‘ D’].
A.4 Further impact scenarios on general conversation.
We shortly discuss about how this method may be applied to general conversational scenarios of LLMs. For processing contexts such as chronological history of conversations, the ordering is important, and the original LLM remains the better choice for this case. However, in subsets of conversational tasks requiring order invariance (e.g., Sets, Tables, or RAG contexts), our method enhances unbiased understanding, as demonstrated mainly in Lost-in-the-Middle benchmark. Here, RoToR achieves a significant 7-9% average accuracy gain over the original LLM, very consistently across all setups (doc indexing and ndoc) for all choices of the ordering algorithm, with lower standard deviation than the original model.
A.5 Selection of $\alpha$ for Selective Routing on MMLU
$\alpha$ = no Selective Routing -0.5 -0.4 -0.3 -0.2 -0.1 0 0.1 0.2 0.3 0.4 0.5
Validation set (1531) <— bias towards invariant model —— bias towards orig model —>
Selective Routing (orig, pine) 64.9 65.3 65.8 66.4 67.3 67.4 67.6 67.5 67.9 67.8 68.0 67.9
Selective Routing (orig, ours-lexical) 63.9 64.1 64.5 65.4 65.8 67.0 67.4 68.1 68.2 68.1 67.9 67.9
Selective Routing (orig, ours-monot5) 66.0 66.1 66.3 66.7 67.0 67.6 68.0 68.1 68.1 68.1 67.7 67.9
Test set (14015)
Selective Routing (orig, pine) 64.8 64.9 65.4 66.0 66.8 67.5 68.4 68.5 68.5 68.4 68.3 68.3
Selective Routing (orig, ours-lexical) 63.2 63.5 64.2 65.2 66.2 67.3 68.0 68.4 68.5 68.5 68.4 68.3
Selective Routing (orig, ours-monot5) 64.2 64.4 64.8 65.5 66.4 67.3 68.1 68.5 68.4 68.5 68.3 68.3
Table 6: Reporting full ablation results on application of Selective Routing. $\alpha$ = 0.2 was the best for the validation set, which was then applied to obtain the reported results for all models.
We report $\alpha$ is a hyperparameter that can be tuned per-dataset. We searched its value in the range of -0.5 to 0.5 with a step size of 0.1 using the validation split of MMLU [^7] https://huggingface.co/datasets/cais/mmlu/viewer/abstract_algebra/validation on RoToR with lexical sorting, and applied the found value (0.2) on the test set to obtain the reported results for all models. We report the full variation of Selective Routing results on the investigated $\alpha$ value at Tab. 6.
A.6 Replication details on the runtime experiment
Apart from the theoretical runtime efficiency, we measured the actual end-to-end runtime in seconds, to better analyze the practical runtime efficiency between PINE and RoToR. The runtime of each experiment was measured on an ASUS ESC8000-E11 server featuring dual 4th Gen Intel Xeon Scalable processors, 64 CPU threads, 1.1 TB of RAM across 32 DIMM slots, and 8 NVIDIA A6000 GPUs with 48 GB of memory each. Except for the experiments on Llama-3.1-70B-Instruct, we only use a single A6000 GPU for all of the experiments.
A.7 Input data examples
To get a better understanding of the input and output format of the lost-in-the middle dataset and structured knowledge grounding dataset, we provide example inputs for each data format.
lost in the middle Prefix:
<|begin_of_text|><|start_header_id|>system<|end_header_id|>
You are a helpful, respectful and honest assistant. Always answer as helpfully as possible, while being safe. Please ensure that your responses are socially unbiased and positive in nature. If a question does not make any sense, or is not factually coherent, explain why instead of answering something not correct. If you don’t know the answer to a question, please don’t share false information.<|eot_id|><|start_header_id|>user<|end_header_id|> Write a high-quality answer for the given question using only the provided search results (some of which might be irrelevant). Parallel texts:
Document [1](Title: List of Nobel laureates in Physics) The first …
Document [10](Title: Nobel Prize in Chemistry) on December 10, the … Suffix:
Question: who got the first nobel prize in physics<|eot_id|><|start_header_id|>assistant
<|end_header_id|>
Figure 7: Example input for the lost in the middle dataset.
lost in the middle no indexing Prefix:
<|begin_of_text|><|start_header_id|>system<|end_header_id|>
You are a helpful, respectful and honest assistant. Always answer as helpfully as possible, while being safe. Please ensure that your responses are socially unbiased and positive in nature. If a question does not make any sense, or is not factually coherent, explain why instead of answering something not correct. If you don’t know the answer to a question, please don’t share false information.<|eot_id|><|start_header_id|>user<|end_header_id|> Write a high-quality answer for the given question using only the provided search results (some of which might be irrelevant). Parallel texts:
[Document Title: List of Nobel laureates in Physics] The first …
[Document Title: Nobel Prize in Chemistry] on December 10, the … Suffix:
Question: who got the first nobel prize in physics<|eot_id|><|start_header_id|>assistant
<|end_header_id|>
Figure 8: Example input for the lost in the middle dataset, without indexing by numbers.
WebQSP Prefix:
<|begin_of_text|><|start_header_id|>system<|end_header_id|>
Below are the facts in the form of the triple meaningful to answer the question. Answer the given question in a JSON format, such as "Answer": "xxx". Only output the JSON, do NOT say any word or explain. <|eot_id|><|start_header_id|>user<|end_header_id|> Parallel texts:
(Ferdinand Magellan, occupation, seafarer)
(Ferdinand Magellan, country of citizenship, Kingdom of Portugal) (Ferdinand Magellan, date of birth, time: +1480-00-00) (Ferdinand Magellan, instance of, human) (Ferdinand Magellan, date of death, time: +1521-04-27) (Ferdinand Magellan, sex or gender, male) (Ferdinand Magellan, killed by, Lapulapu) Suffix:
Question: what country did magellan represent?, Answer: <|eot_id|><|start_header_id|>assistant <|end_header_id|>
Gold Answer(s):
(‘Reino de Portugal’, ‘Portuguese Kingdom’, ‘Kingdom of Portugal’)
Example generated output:
{"Answer": "Portugal"} (Parsed to: Portugal)
Figure 9: Example input for the WebQSP dataset.
Mintaka Prefix:
<|begin_of_text|><|start_header_id|>system<|end_header_id|>
Below are the facts in the form of the triple meaningful to answer the question. Answer the given question in a JSON format, such as "Answer": "xxx". Only output the JSON, do NOT say any word or explain. <|eot_id|><|start_header_id|>user<|end_header_id|> Parallel texts:
(Super Bowl XLII, winner, New York Giants)
(Super Bowl XLII, participating team, New York Giants) (Super Bowl XLII, point in time, time: +2008-02-03) (Super Bowl XLII, followed by, Super Bowl XLIII) (Super Bowl XLII, location, State Farm Stadium) (Super Bowl XLII, sport, American football) (Super Bowl XLII, instance of, Super Bowl) Suffix:
Question: which team did the super bowl xlii mvp play for?, Answer: <|eot_id|><|start_header_id|> assistant <|end_header_id|>
Gold Answer(s):
(‘NYG’, ‘Giants’, ‘NY Giants’, ‘New York Giants’)
Example generated output:
{"Answer": "New York Giants"} (Parsed to: New York Giants)
Figure 10: Example input for the Mintaka dataset.
Appendix B Why is removing index bias an important detail for invariant models to be effective?
The alphabetic index (A/B/C/D) introduced in Fig. 1 associated with each segment, reportedly introduces token bias Wei et al. (2024) of preferring the choice marked as ‘A.’ The same thing can be applied to listwise inputs with simple numeric indexing (1/2/3/4), which was the case for the lost-in-the middle benchmark. While a standard model with no modifications on positional encoding correctly places contexts indexed A before contexts indexed with D by positional encoding, an invariant model sees contexts in an order-agnostic way, meaning that the alphabetical indexing may not always be interpreted sequentially and thus can confuse the model from accurately interpreting the contexts. For example, even for cases where the index ordering of the input was in alphabetical order (A->B->C->D), the ordering-invariant model may interpret contexts with (C->A->B->D) at one point (e.g., when the query is D on self attention), which can cause unnatural, out-of-distribution representation, leading to decreased performance.
Appendix C Statisticial significance before and after shuffling segments
We have calculated the paired one-tailed t-tests performed on both the baseline (“original”) model and our proposed method (RoToR) at Tab. 8 using the results at Tab. 7 on whether the performance differences between initial ordering v.s. shuffled ordering are statistically significant on the original model and ours. Note that the performance on the Lost-in-the Middle (LitM) was not reported since the LitM benchmark didn’t provide initial ordering information. Specifically, we test whether the mean performance difference, (Before Shuffle - After Shuffle), deviates significantly from zero. As shown in Table 8, the results indicate that while the original model does exhibit a statistically significant drop in performance upon shuffling, our RoToR method does not, suggesting increased robustness to segment-order perturbations.  [^8] The calculation were validated using the online t-test calculator. https://www.mathportal.org/calculators/statistics-calculator/t-test-calculator.php
Original Model RoToR-lexical
Before Shuff. After Shuff. Diff. Before Shuff. After Shuff. Diff.
MMLU 68.3 65.5 2.8 68.5 65.7 2.8
WebQSP, ndoc 30 57.6 57.0 0.6 58.8 58.7 0.1
WebQSP, ndoc 50 57.5 56.5 1.0 58.8 58.7 0.1
Mintaka, ndoc 30 48.7 47.9 0.8 51.2 50.9 0.3
Mintaka, ndoc 50 48.6 48.1 0.5 50.8 50.7 0.1
Table 7: Performance of the Original model and RoToR before and after shuffling.
Original RoToR
Degrees of Freedom 4
Mean Difference 1.14 0.68
t-Statistic 3.01 1.43
Critical Value 2.132
Statistically Significant Yes No
Table 8: Paired one-tailed t-test results comparing the original model and ours.
Derivation for the original model.
Let the five paired differences (Before -- After) be $\{d_{1},d_{2},d_{3},d_{4},d_{5}\}$. Mean Difference: $\bar{d}=\tfrac{1}{5}\sum_{i=1}^{5}d_{i}$. In this case, $\bar{d}\approx 1.14\%.$ Sample Standard Deviation: $s_{d}=\sqrt{\frac{1}{n-1}\sum_{i=1}^{n}(d_{i}-\bar{d})^{2}}\approx 0.8476.$ Standard Error (SE): $\mathrm{SE}=\tfrac{s_{d}}{\sqrt{n}}\approx 0.379.$ t-Statistic: $t=\tfrac{\bar{d}}{\mathrm{SE}}\approx 3.01,\ (df=4).$ Since the critical value at $df=4$ and $\alpha=0.05$ is $2.132,$ we have $3.01>2.132$ Therefore, the difference is statistically significant.
Derivation for the RoToR model.
Under the same procedure, $\bar{d}\approx 0.68\%,\ s_{d}\approx 1.0628,\ \mathrm{SE}\approx 0.475.$ t-Statistic: $t\approx 1.4307.$ Since $1.4307<2.132,$ there is no significant difference in performance before and after shuffling for RoToR.
Conclusion.
While the original model shows a statistically significant performance drop with shuffled inputs, RoToR remains unaffected, demonstrating greater robustness to segment-order perturbations.

### v2

@@
 In this work, as a first contribution, we identify and overcome two limitations to make zero-shot invariant LMs more practical: (1) training and inference distribution mismatch arising from modifying positional ID assignments to enforce invariance, and (2) failure to adapt to a mixture of order-invariant and sensitive inputs in practical listwise problems.
-To overcome, we propose (1) RoToR, a zero-shot invariant LM for genuinely order-invariant inputs with minimal modifications of positional IDs, and (2) Selective Routing, an adaptive framework that handles both order-invariant and order-sensitive inputs in listwise tasks.
-On the Lost in the middle (LitM), Knowledge Graph Question Answering (KGQA), and MMLU benchmarks, we show that RoToR with Selective Routing can effectively handle practical listwise input tasks in a zero-shot manner.
+Then, to overcome these issues we propose (1) RoToR, a zero-shot invariant LM for genuinely order-invariant inputs with minimal modifications of positional IDs, and (2) Selective Routing, an adaptive framework that handles both order-invariant and order-sensitive inputs in listwise tasks.
+On the Lost in the middle (LitM), Knowledge Graph QA (KGQA), and MMLU benchmarks, we show that RoToR with Selective Routing can effectively handle practical listwise input tasks in a zero-shot manner.
 [^1] https://github.com/soyoung97/RoToR
@@
 For example, for inputs such as sets, tables, databases, or multiple-choice questions, the ordering of the input segments—e.g., rows in a table or elements in an unordered set—require an order-agnostic understanding.
-We refer to such inputs as “order-invariant inputs,” which LMs reportedly struggle on.
-For example, in LLM-as-a judge scenarios, they exhibit a preference of up to 75% for the first answer in pairwise inputs Zheng et al.
+We refer to such inputs as “order-invariant inputs,” on which LMs reportedly struggle.
+For example, in LLM-as-a-judge scenarios, LMs exhibit a preference of up to 75% for the first answer in pairwise inputs Zheng et al.
 (2024b), and ranking between LMs can change up to 8 positions in different orderings of multiple choice questions on MMLU Alzahrani et al.
 (2024).
-Such results question the reliability of LM outputs on order-invariant inputs.
+Such results question the reliability of LMs on order-invariant inputs.
 Meanwhile, existing methods for enforcing invariance to LMs showed limited effectiveness in real-world tasks, which we hypothesize to arise from the following limitations.
 First, training and inference distribution mismatch due to the positional ID re-assignment of zero-shot order-invariant LMs: Fig.
-1 illustrates how self-attention is altered in there models.
+1 illustrates how self-attention is altered in these models.
 Unlike the original non-invariant model which always assigns position IDs in a causal, ascending manner, order-invariant models either eliminate inter-segment attention, such as PCW Ratner et al.
@@
 (2024) in Fig.
-1b, re-ordering segments using an importance score, placing similar segments closer to the query.
+1b, re-ordering segments using pairwise similarity, placing similar segments closer to the query.
 For each query segment, it computes segment-wise query-key attention (for each attention head in each decoder layer) and re-assigns position IDs of segments as keys.
-This query-dependent segment ordering leads to excessively frequent alterations on positional ID assignments.
-Frequent re-assignments can also confuse the model and risk collisions which violates the invariance property (e.g., multiple segments having the same importance score).
+This query-dependent segment ordering leads to excessively frequent alterations of positional ID assignments.
+Frequent re-assignments can also confuse the model and risk collisions which violate the invariance property (e.g., multiple key segments having the same similarity to a query).
 To overcome this, we propose a query-agnostic global sorting with circular arrangement for order-invariant positional ID assignment.
@@
 1c, where RoToR only needs a single global ordering (e.g., A->B->O->K->G) with no extra attention computation.
-The ordering of segments on suffix tokens remain in a fixed order, since it does not rely on their similarity to the query.
+The ordering of segments on suffix tokens remains in a fixed order, since it does not rely on their similarity to the query.
 Finally, we propose three different global sorting algorithms for RoToR, and demonstrate that they consistently outperform previous order-invariant models.
-Second, in practical listwise inputs, order-invariant task may include order-sensitive inputs which requires order-specific understanding.
+Second, for practical listwise inputs, order-invariant tasks may partially include order-sensitive inputs that require order-specific understanding.
 For example, the (d) None of the above option in MMLU cannot be reordered.
-For such a “mixed” nature requires to handle each of them adaptively by a simple Selective Routing method.
-Selective Routing adapts to input by routing between two models, invariant and non-invariant (original), based on their confidence scores of predictions.
-Experiments on the MMLU benchmark show that Selective Routing effectively handles datasets with order-invariant and sensitive inputs, and achieve better order robustness while maintaining the original performance.
+Such a “mixed” nature requires handling each of the cases adaptively, for which we propose a simple Selective Routing method.
+Selective Routing adapts to a given input by routing between two models, invariant and non-invariant (original), based on the confidence scores of their predictions.
+Experiments on the MMLU benchmark show that Selective Routing effectively handles datasets with order-invariant and sensitive inputs, and achieves better order robustness while maintaining the original performance.
 In summary, our contributions are as follows: 1.
@@
 (2024); Cai et al.
-(2023) models, leaving decoder-only models, which account for most of the current LLMs, for more consideration.
+(2023) models, they leave decoder-only models, which account for most of the current LLMs, for more consideration.
 In contrast to transformer encoders that use bidirectional attention which is invariant by nature Lee et al.
-(2019), transformer decoders use causal attention in order to learn causal relation signals, which is not invariant by nature Haviv et al.
+(2019), transformer decoders use causal attention to learn causal relation signals, which is not invariant by nature Haviv et al.
 (2022a).
@@
 (2024); Wang et al.
-(2024).
+(2024) and is harder to mitigate.
 2.2 Zero-shot order-invariance for LLMs
@@
 3.1 Baseline: Order-invariant causal LMs
-In this section, we briefly overview the existing work on turning decoder-only LLM architectures into order-invariant models by adjusting attention mechanism, and review their limitations.
+In this section, we briefly overview the existing work on endowing decoder-only models on order-invariance by adjusting attention mechanism, and review their limitations.
 Isolated parallel processing
@@
 PINE dynamically re-assigns positional IDs depending on a token’s role as a query or a key, instead of using a single fixed positional ID.
-For Case 1: When “ana” (T4) in segment S2 acts as a query, S2 is assigned the largest position IDs (5) so it can attend to all previous segments.
+For Case 1: When “ana” (T4, S2) acts as a query, S2 is assigned the largest position IDs (5) so it can attend to all previous segments.
 Within S2, the tokens “ban” (4) and “ana” (5) still maintain their local causal order (i.e., smaller to larger positions).
@@
 Total ndoc (segments) 10 20 30
-Gold idx at: 0 4 9 avg. 0 4 9 14 19 avg. 0 4 9 14 19 24 29 avg.
-Indexing bias present
-Original 52.8 51.1 46.7 50.2 55.0 50.7 50.5 50.2 47.5 50.7 56.1 50.8 51.5 51.1 50.8 50.5 48.9 51.4
-PCW 12.0 11.9 12.1 12.0 03.4 03.7 03.8 03.9 03.6 05.1 02.1 02.1 02.0 01.9 02.0 02.2 02.0 02.0
-Set-Based Prompting 40.8 40.7 40.8 40.8 25.6 25.8 25.7 25.5 25.3 25.6 15.8 15.9 16.1 16.0 16.1 15.7 15.8 15.9
-PINE 59.2 56.8 57.7 57.9 56.2 55.1 54.5 54.2 55.3 55.5 56.1 53.6 53.3 53.9 53.2 53.7 54.5 54.0
-RoToR-lexical 63.6 60.5 59.1 61.1 52.6 58.6 57.8 56.8 58.6 57.6 64.6 58.9 56.9 56.2 57.1 56.2 57.1 58.1
-RoToR-reversed lexical 61.5 60.8 60.6 61.0 60.8 58.6 59.7 60.5 59.3 60.0 61.1 57.5 58.2 58.6 58.3 59.3 59.2 58.9
-RoToR-reranking 61.4 61.4 61.7 61.5 62.3 59.2 59.1 59.4 59.8 60.2 62.2 58.6 58.4 58.6 58.3 58.5 59.8 59.2
-RoToR-freq 62.8 61.1 59.5 61.1 62.9 58.8 56.7 57.4 58.0 59.1 61.7 58.2 56.9 56.1 56.4 55.4 56.8 57.4
-Indexing bias removed
-Original 54.7 53.0 50.2 52.6 54.8 52.6 52.8 52.4 51.0 52.7 55.6 51.5 52.4 52.8 52.1 52.3 53.0 52.8
-PCW 12.4 11.9 12.2 12.2 03.7 04.0 04.0 04.0 03.9 03.9 02.3 01.8 02.0 02.0 02.1 02.0 02.0 02.0
-Set-Based Prompting 42.5 42.5 42.5 42.5 26.3 26.3 26.3 26.3 26.3 26.3 14.1 14.1 14.1 14.1 14.1 14.1 14.1 14.1
-PINE 58.6 58.8 59.0 58.8 56.2 55.7 55.5 55.7 55.5 55.7 54.2 54.8 54.3 53.7 54.8 54.2 54.0 54.3
-RoToR-lexical 61.4 61.6 61.6 61.5 61.4 59.8 59.6 59.6 59.8 60.0 59.2 59.5 59.4 59.1 59.0 59.3 59.1 59.2
-RoToR-reversed lexical 61.6 61.8 61.8 61.8 58.9 59.3 58.8 58.6 58.7 58.8 57.9 58.2 57.9 57.4 57.9 57.6 57.5 57.8
-RoToR-reranking 61.2 61.4 61.2 61.3 60.9 61.0 61.2 61.2 61.2 61.1 60.9 60.7 60.7 60.7 60.8 60.8 60.7 60.8
-RoToR-freq 61.0 61.1 61.1 61.1 60.4 60.3 58.6 60.2 60.0 59.9 59.3 60.4 59.7 59.5 59.5 59.6 59.2 59.6
+Gold idx at: 0 4 9 0 4 9 14 19 0 4 9 14 19 24 29
+Llama-3.1-8B-Instruct
+Original 54.7 53.0 50.2 54.8 52.6 52.8 52.4 51.0 55.6 51.5 52.4 52.8 52.1 52.3 53.0
+PCW 12.4 11.9 12.2 3.7 4.0 4.0 4.0 3.9 2.3 1.8 2.0 2.0 2.1 2.0 2.0
+Set-Based Prompting 42.5 42.5 42.5 26.3 26.3 26.3 26.3 26.3 14.1 14.1 14.1 14.1 14.1 14.1 14.1
+PINE 58.6 58.8 59.0 56.2 55.7 55.5 55.7 55.5 54.2 54.8 54.3 53.7 54.8 54.2 54.0
+RoToR-lexical 61.4 61.6 61.6 61.4 59.8 59.6 59.6 59.8 59.2 59.5 59.4 59.1 59.0 59.3 59.1
+RoToR-reversed lexical 61.6 61.8 61.8 58.9 59.3 58.8 58.6 58.7 57.9 58.2 57.9 57.4 57.9 57.6 57.5
+RoToR-MonoT5 61.2 61.4 61.2 60.9 61.0 61.2 61.2 61.2 60.9 60.7 60.7 60.7 60.8 60.8 60.7
+RoToR-Freq. 61.0 61.1 61.1 60.4 60.3 58.6 60.2 60.0 59.3 60.4 59.7 59.5 59.5 59.6 59.2
+Qwen1.5-4B-Chat
+Original 61.3 54.8 53.1 59.5 49.1 47.9 45.9 48.3 56.8 45.6 44.9 44.6 45.3 43.5 48.3
+PINE 57.2 57.4 57.0 48.6 48.2 48.2 48.1 48.9 46.4 45.9 46.7 46.6 46.4 46.4 46.3
+RoToR 58.5 58.4 58.1 49.9 49.7 49.6 49.8 49.9 44.6 44.8 44.7 44.7 44.9 44.8 44.7
+RoToR-MonoT5 58.9 58.5 58.7 52.2 52.1 52.1 52.2 52.6 50.6 50.7 50.5 50.6 50.5 50.6 50.4
+RoToR-Freq. 56.7 56.9 56.9 51.9 51.5 51.8 51.6 52.4 46.8 46.7 46.7 46.4 47.0 46.8 46.6
+Qwen1.5-7B-Chat
+Original 72.5 63.3 62.9 72.5 58.5 56.1 56.0 58.2 73.1 58.6 55.8 53.3 53.2 52.5 57.5
+PINE 65.4 65.5 66.3 59.1 59.4 59.1 58.6 59.2 58.0 55.3 55.7 56.3 55.1 55.8 56.1
+RoToR 68.6 68.7 68.6 62.6 62.9 62.7 63.0 62.7 57.0 57.3 59.7 57.4 57.3 62.8 57.0
+RoToR-MonoT5 68.8 69.4 69.0 65.2 65.5 65.0 64.9 65.0 62.6 62.8 62.9 62.7 62.9 62.8 62.5
+RoToR-Freq. 68.2 68.4 68.4 62.6 62.9 62.8 62.7 62.3 59.5 59.8 59.7 59.6 59.7 59.7 59.7
 […]
-WebQSP (Count = 1466) Mintaka (Count = 4000)
-N=30 N=50 N=30 N=50
-Acc EM F1 Avg. Acc EM F1 Avg. Acc EM F1 Avg. Acc EM F1 Avg.
-Initial (no shuffling of segments)
-Orig. 71.3 43.2 58.3 57.6 71.6 43.0 58.1 57.5 50.2 44.0 51.9 48.7 50.0 44.0 51.7 48.6
-PINE 72.0 44.1 59.0 58.4 72.4 44.4 58.5 58.4 51.5 45.0 52.6 49.7 51.6 45.1 52.6 49.8
-RoToR 73.7 43.4 59.2 58.8 73.9 43.5 59.0 58.8 53.1 46.5 54.1 51.2 52.9 46.0 53.6 50.8
-RoToR (MonoT5) 72.9 43.6 59.2 58.6 73.9 43.9 59.1 59.0 51.6 45.0 52.5 49.7 52.4 45.4 52.8 50.2
-RoToR (Freq) 73.5 43.5 59.3 58.8 74.1 43.9 59.3 59.1 52.6 46.1 53.7 50.8 53.1 46.4 53.7 51.0
+Llama-3.1-8B-Instruct Qwen1.5-4B-Chat Qwen1.5-7B-Chat
+N = 30 N = 50 N = 30 N = 50 N = 30 N = 50
+Method Acc. EM F1 Acc. EM F1 Acc. EM F1 Acc. EM F1 Acc. EM F1 Acc. EM F1
+Initial, no shuffling of segments
+Original 50.2 44.0 51.9 50.0 44.0 51.7 30.7 27.9 34.9 31.6 28.6 35.8 31.5 27.8 35.4 31.7 28.0 35.7
+PINE 51.5 45.0 52.6 51.6 45.1 52.6 31.6 28.7 35.6 31.6 28.8 35.3 32.3 28.8 36.4 32.0 28.5 35.9
+RoToR 53.1 46.5 54.1 52.9 46.0 53.6 32.0 29.0 35.7 32.7 29.6 36.2 34.3 29.8 37.7 34.3 30.1 38.0
+RoToR-MonoT5 51.6 45.0 52.5 52.4 45.4 52.8 32.3 29.1 36.2 32.3 29.3 35.9 32.9 28.4 36.3 32.9 28.9 36.6
+RoToR-Freq. 52.6 46.1 53.7 53.1 46.4 53.7 32.3 29.2 36.0 32.3 29.2 35.9 33.7 29.5 37.2 33.5 29.5 37.2
 After shuffling segments, averaged over 3 seeds
-Orig. 70.9 42.4 57.8 57.0 $\pm$ 0.5 70.7 41.6 57.2 56.5 $\pm$ 0.2 49.5 43.3 51.0 47.9 $\pm$ 0.0 49.7 43.5 51.0 48.1 $\pm$ 0.4
-PINE 71.9 43.9 59.1 58.1 $\pm$ 0.0 72.5 44.6 58.5 58.5 $\pm$ 0.3 51.8 45.2 52.8 50.0 $\pm$ 0.1 51.8 45.3 52.7 49.9 $\pm$ 0.2
-RoToR 73.6 43.2 59.1 58.7 $\pm$ 0.1 74.0 43.3 58.8 58.7 $\pm$ 0.1 52.8 46.2 53.8 50.9 $\pm$ 0.0 52.7 45.9 53.5 50.7 $\pm$ 0.0
-RoToR (MonoT5) 72.6 43.4 59.2 58.3 $\pm$ 0.1 73.9 43.7 58.9 58.8 $\pm$ 0.1 51.6 45.0 52.6 49.7 $\pm$ 0.1 52.2 45.2 52.8 50.1 $\pm$ 0.2
-RoToR (Freq) 72.9 43.2 59.0 58.4 $\pm$ 0.0 73.7 43.8 59.1 58.9 $\pm$ 0.1 52.5 45.9 53.5 50.6 $\pm$ 0.1 53.1 46.4 53.7 51.1 $\pm$ 0.0
-Table 2: Knowledge Graph Question Answering (KGQA) results, with standard deviation of the average scores with $\pm$. N refers to number of top-k segments per query.
-Initial
-ordering
-Reversed
-ordering
-Avg. of all possible
-(4!-1) re-orderings
-Baseline: Single use of order-invariant models
-Orig. 68.3 64.8 65.5 $\pm$ 1.0
+Original 49.5 43.3 51.0 49.7 43.5 51.0 30.1 27.5 34.7 30.3 27.6 35.0 31.4 27.3 35.0 31.6 27.9 35.5
+$\hookrightarrow$ stdev. ($\pm$) 0.07 0.14 0.17 0.34 0.28 0.46 0.41 0.34 0.43 0.26 0.24 0.35 0.26 0.28 0.29 0.40 0.56 0.42
+PINE 51.8 45.2 52.8 51.8 45.3 52.7 31.5 28.7 35.6 31.5 28.7 35.3 32.3 28.8 35.7 31.7 28.2 35.7
+$\hookrightarrow$ stdev. ($\pm$) 0.05 0.07 0.16 0.15 0.16 0.19 0.20 0.18 0.13 0.17 0.20 0.21 0.17 0.20 0.13 0.18 0.16 0.14
+RoToR 52.8 46.2 53.8 52.7 45.9 53.5 31.8 28.8 35.5 32.5 29.6 36.1 34.2 29.9 37.7 34.2 30.1 38.0
+$\hookrightarrow$ stdev. ($\pm$) 0.05 0.05 0.02 0.05 0.09 0.04 0.05 0.02 0.09 0.11 0.06 0.09 0.09 0.07 0.06 0.06 0.05 0.04
+RoToR-MonoT5 51.6 45.0 52.6 52.2 45.2 52.8 32.4 29.2 36.3 32.3 29.4 35.9 33.0 28.8 36.5 32.8 28.8 36.5
+$\hookrightarrow$ stdev. ($\pm$) 0.12 0.06 0.10 0.16 0.18 0.18 0.04 0.02 0.13 0.16 0.13 0.07 0.12 0.09 0.07 0.16 0.09 0.07
+RoToR-Freq. 52.5 45.9 53.5 53.1 46.4 53.7 32.3 29.3 36.0 32.4 29.3 36.1 33.8 29.6 37.4 33.7 29.6 37.4
+$\hookrightarrow$ stdev. ($\pm$) 0.10 0.15 0.11 0.02 0.07 0.03 0.13 0.16 0.09 0.09 0.04 0.06 0.04 0.00 0.09 0.04 0.16 0.22
+Table 2: Results on the Mintaka (KGQA) dataset on different models, with standard deviation of the average scores with $\pm$. N refers to number of top-k segments per query.
+Llama-3.1-8B-Instruct Qwen1.5-4B-Chat Qwen1.5-7B-Chat
+Method Init. Rev. Avg. Init. Rev. Avg. Init. Rev. Avg.
+Orig. 68.3 64.8 65.5 $\pm$ 1.0 53.6 51.9 52.6 $\pm$ 0.6 60.1 56.6 58.6 $\pm$ 0.9
 PCW 57.0 55.1 56.1 $\pm$ 1.1
 Set-Based Prompting 31.1 33.0 31.6 $\pm$ 0.8
-PINE 64.8 63.3 63.6 $\pm$ 0.7
-RoToR 63.2 62.6 62.8 $\pm$ 0.7
-RoToR (MonoT5) 64.2 62.9 63.5 $\pm$ 0.5
-RoToR (Freq) 64.3 63.5 63.6 $\pm$ 0.6
-With Selective Routing
-Orig. 68.3 64.8 65.5 $\pm$ 1.0
-Orig + PINE 68.5 65.1 65.7 $\pm$ 0.9
-Orig + RoToR 68.5 65.0 65.7 $\pm$ 0.9
-Orig + RoToR (MonoT5) 68.4 65.2 65.8 $\pm$ 0.9
-Orig + RoToR (Freq) 68.5 65.3 65.8 $\pm$ 0.8
-Orig + RoToR (Optimal) 75.0 71.9 72.7 $\pm$ 1.0
-Table 3: Improving applicability to general listwise tasks (MMLU, N=4) with Selective Routing, which includes both order-invariant and order-sensitive examples.
-RoToR with Selective Routing shows improved performance and stability across re-orderings of input, and its high upper-bound (Optimal) implies the potential for further improvements.
+PINE 64.8 63.3 63.6 $\pm$ 0.7 50.5 49.3 49.4 $\pm$ 0.5 57.0 54.4 55.8 $\pm$ 0.9
+RoToR 63.2 62.6 62.8 $\pm$ 0.7 49.6 47.7 48.3 $\pm$ 0.7 56.5 55.8 56.2 $\pm$ 0.6
+$\hookrightarrow$ + S.R. 68.5 65.1 65.7 $\pm$ 0.9 53.7 51.8 52.6 $\pm$ 0.6 60.1 57.4 58.8 $\pm$ 0.7
+RoToR - MonoT5 64.2 62.9 63.5 $\pm$ 0.5 49.7 47.6 48.7 $\pm$ 0.7 56.2 54.4 55.5 $\pm$ 0.7
+$\hookrightarrow$ + S.R. 68.4 65.2 65.8 $\pm$ 0.9 53.8 51.9 52.6 $\pm$ 0.6 60.1 57.3 58.7 $\pm$ 0.8
+RoToR - Freq. 64.3 63.6 63.8 $\pm$ 0.6 49.9 47.6 48.7 $\pm$ 0.5 56.4 54.7 55.7 $\pm$ 0.7
+$\hookrightarrow$ + S.R. 68.5 65.3 65.8 $\pm$ 0.8 53.7 52.3 52.6 $\pm$ 0.6 60.0 57.3 58.6 $\pm$ 0.8
+RoToR + S.R. (Optim.) 75.0 71.9 72.7 $\pm$ 1.0 61.8 60.1 61.1 $\pm$ 1.0 68.1 66.2 67.2 $\pm$ 0.7
+Table 3: Improving applicability to general listwise tasks (MMLU, N=4) with Selective Routing (S.R), which includes both order-invariant and order-sensitive examples.
+RoToR with Selective Routing shows improved performance and stability across re-orderings of input, and its high upper-bound (Optim.) implies the potential for further improvements.
 4.1 Baselines
@@
 (2024).
-We use the LLaMA 3.1 AI (2024) 8B-Instruct444meta-llama/Meta-Llama-3.1-8B-Instruct as our backbone model for all of our experiments [^5] Results on 70B model is available in Appendix, Tab.
-5..
-As our method doesn’t need training, a single A6000 GPU was sufficient to run all of the experiments with our primary backbone model.
+We use the LLaMA 3.1 AI (2024) 8B-Instruct444meta-llama/Meta-Llama-3.1-8B-Instruct, Qwen1.5-4B-Chat and Qwen1.5-7B-Chat [^5] Qwen/Qwen1.5-4/7B-Chat as our backbone model for all of our experiments.
+As our method doesn’t need training, a single A6000 GPU was sufficient to run all of the experiments except for the Llama-3.1-70B-Instruct model (Appendix, Tab.
+6).
+We also conduct experiments on a subset of benchmarks (LitM and MMLU) on the runtime latency, perplexity, and collision rate of PINE and RoToR, to further validate our claims on Sec.
+3.2.
 4.2 Benchmarks with listwise inputs
-We experiment with three benchmarks involving real-world listwise input data. Examples of exact inputs and outputs are provided in Appendix A.7. All reported scores are rounded to the nearest tenth.
-Knowledge Graph Question Answering (KGQA)
-In KGQA tasks, the model takes facts over knowledge graphs represented as (subject, relation, object), and answers the given question based on the given facts.
+We experiment with three benchmarks involving real-world listwise input data.
+Examples of exact inputs and outputs are provided in Appendix A.7.
+All reported scores are rounded to the nearest tenth, except for the standard deviation (rounded to the second decimal place).
+Knowledge Graph QA (KGQA) In KGQA tasks, the model takes facts over knowledge graphs represented as (subject, relation, object), and answers the given question based on the given facts.
 We basically follow the KGQA dataset preprocessing and evaluation setup from  Baek et al.
-(2023), which uses WebQSP Yih et al.
-(2016) modified from  Berant et al.
-(2013) and Mintaka Sen et al.
+(2023), which uses Mintaka Sen et al.
 (2022) with Wikidata for knowledge source, and use the Exact Match (EM), Accuracy (Acc), and F1 score metric for evaluation.
@@
 Replication details and example dataset format are at Appendix Sec.
-A.3, Fig.
-9 and Fig.
-10, respectively.
+A.3 and Fig.
+10.
 Along with measuring the performance of the initial input ordering, we report performance after we shuffle the order of segments with 3 different seeds to see shuffle robustness.
@@
 (2024), which draws from 2655 queries in the Natural Questions (NQ) dataset.
-It provides sets of (10, 20, 30) passages, placing the gold passage at predetermined positions (e.g., 0, 4, 9) and filling the remaining slots with irrelevant passages, testing the model’s robustness to passage ordering.
+It provides sets of (10, 20, 30) passages, placing the gold passage at predetermined positions (e.g., 0, 4, 9) and filling the remaining slots with irrelevant passages.
 Following  Liu et al.
-(2024), we measure answer quality with the best_subspan_em metric.
-Experiments on LitM found that eliminating the effect of index bias is another important detail for invariant models to be effective: (Appendix Sec.
-B) thus, we also experiment with a variant of the benchmark with index bias eliminated (prompt at Appendix Fig.
-8) in order to further investigate the impact with and without index bias.
+(2024), the best_subspan_em metric is used.
+Experiments on LitM found that eliminating the effect of index bias is another important detail for measuring true order robustness: (Appendix Sec.
+B).
+Thus, we report experiments with index bias eliminated.
+The exact prompt and full results including index bias is reported at Appendix Fig.
+8 and Sec.
+A.1.
 MMLU The Massive Multitask Language Understanding (MMLU) benchmark Hendrycks et al.
-(2021) consists of 57 diverse sub-tasks with a total of 14,015 queries to measure general performance of LMs about the knowledge of the world.
+(2021) (prompts at Appendix Fig.
+11) consists of 57 diverse sub-tasks with a total of 14,015 queries to measure general performance of LMs about the knowledge of the world.
 Despite its popularity, many works report that performance fluctuates heavily depending on the order of choices Gupta et al.
@@
 (2024a) and is widely investigated to measure the positional bias of the model.
-On experiments on MMLU, we notice that a lot of proportions consist of ordering-sensitive inputs, which showed the effectiveness of adaptively applying Selective Routing.
-In addition to measuring the performance with the original ordering, we report the average performance for all possible (4!-1) re-orderings.
-4.3 Further comparison with PINE
-We additionally conduct experiments on a subset of benchmarks (LitM and MMLU) on the runtime latency, perplexity, and collision rate of PINE and RoToR, to validate our claims on Sec. 3.2.
+We notice that a lot of proportions consist of ordering-sensitive inputs, which showed the effectiveness of adaptively applying Selective Routing.
+We additionally report the average performance for all possible (4!-1) re-orderings.
 5 Results & Analysis
-Overview
-We show results for KGQA in Tab. 2, and results for MMLU in Tab. 3. Results for LitM are in Tab. 1, with a visualization in Appendix Fig. 5. We use lexical sorting for RoToR unless stated otherwise.
-Impact of removing index bias on LitM
-On LitM, results in Tab.
-1 show that, while invariant LMs show stable performances regardless of the gold index when index bias is removed (as described in Sec.
-4.2) (Appendix Fig.
-5(b)), fluctuations are observed if index bias is present (Appendix Fig.
-5(a)).
-This implies that index bias serves as an implicit source of additional positional bias, and invariant LMs benefit from removing it.
+We report results for KGQA in Tab. 2, and results for MMLU in Tab. 3. Results for LitM are in Tab. 1, with a visualization in Appendix Fig. 5. We use lexical sorting for RoToR unless stated otherwise.
 Effectiveness of RoToR
-We observe that shuffling input segments leads to non-trivial performance degradations in the original model, which exhibits a statistically significant performance drop on our experimented dataset (one-tailed t-test, p < 0.05).
-In contrast, our proposed RoToR model does not show a statistically significant difference in performance before and after shuffling, indicating that it is more robust against such perturbations (Appendix C).
+We observe that shuffling input segments leads to non-trivial performance degradations in the original model, which exhibits a statistically significant performance drop on our experimented dataset (two-tailed t-test, p < 0.05, Appendix C).
+In contrast, our proposed RoToR model does not show a statistically significant difference in performance before and after shuffling, indicating that it is more robust against such perturbations.
 On LitM (Tab.
-1), we notice PCW and Set-Based Prompting has impractical performance, with PINE degrading heavily (-4.5%) as number of documents ($k$) increases, while RoToR is less affected.
+1), we notice PCW and Set-Based Prompting has impractical performance, with PINE degrading heavily as number of documents ($k$) increases, while RoToR is less affected.
 On KGQA (Tab.
-2), we show RoToR outperform PINE with lower standard deviation across shuffled segments.
-The effectiveness increases as N increases (30 -> 50).
-Practical scalability with increasing $k$ is critical, but we find that previous order-invariant LMs struggle handling larger $k$ (on KGQA and LitM).
-In contrast, RoToR shows better performance with improved efficiency and robustness.
+2), we show RoToR outperform PINE with lower standard deviation across shuffled segments, consistent with different model architectures.
 Improvements from PINE
@@
 Experiments against comparing RoToR with PINE (Tab.
-4) analyze the following: Runtime: Actual inference times (Appendix Sec.
-A.6) and find that RoToR outperforms PINE substantially, with efficiency gains increasing alongside $n$.
+4) analyze the following: Runtime, scalability: Actual inference times (Appendix Sec.
+A.6) find that RoToR outperforms PINE substantially, with efficiency gains increasing alongside $n$.
 For instance, on LitM (30 docs), RoToR achieves a 43% reduction in total runtime.
+Practical scalability with increasing $k$ is critical, but we find that previous order-invariant LMs struggle handling larger $k$ (on KGQA and LitM).
+In contrast, RoToR shows better performance with improved efficiency and robustness.
 Perplexity: Lower generation perplexity indicates input representations are closer to in-distribution.
@@
 Selective Routing
-Selective Routing on MMLU (Tab.
+MMLU (Tab.
 3) is a representative of a task that involves not only order-invariant, but also order-sensitive (e.g., "None of the above"), inputs.
-Therefore, single use of order-invariant models does not always outperform the original model, limiting applicability of order-invariant models to practical listwise tasks.
-For example, we observe significant performance drops for Set-based Prompting in MMLU, falling short of half the performance of the original model on initial ordering.
-Using RoToR with Selective Routing to handle order-sensitive inputs can be a solution to this case, which outperforms the original model in all possible orderings of candidate choices.
+Therefore, single use of order-invariant models does not always outperform the original model, limiting applicability of order-invariant models to practical listwise tasks, i.e., we observe significant performance drops for Set-based Prompting in MMLU, falling short of half the performance of the original model on initial ordering.
+However, using RoToR with Selective Routing to handle order-sensitive inputs outperforms, or is at least competitive as the original model in all possible orderings of candidate choices.
 Selective Routing improves the generalizability and extends the applicability on practical listwise tasks by adaptively handling order-sensitive inputs.
-We additionally report the the RoToR + Selective Routing (Optimal) performance on Tab.
-3, which improves by a significant 7% over using a single model.
-It was evaluated using a relaxed accuracy metric based on the union of predictions from the original and the invariant (RoToR-lexical) model.
-This highlights the potential of Selective Routing for further accuracy gains through optimizing design choices on routing methods, which we plan to explore in future work.
+The RoToR + Selective Routing (Optimal) performance on Tab.
+3 was evaluated using a relaxed accuracy metric based on the union of predictions from the original and the invariant (RoToR-lexical) model.
+This improves significantly, which highlights the potential of Selective Routing for further accuracy gains through optimizing design choices on routing methods, which we plan to explore in future work.
 Impact of global ordering algorithm
@@
 Our work addresses order-invariance in listwise inputs by identifying core issues in distribution mismatch and adaptive handling of mixed inputs.
-Our proposed RoToR provides a stable zero-shot order-invariant solution that reduces the complexity of positional ID modification, while MoV adaptively routes between invariant and sensitive LMs to handle real-world scenarios.
-Together, these methods demonstrate improved performance and reliability on KGQA, LitM, and MMLU benchmarks.
+Our proposed RoToR provides a stable zero-shot order-invariant solution that reduces the complexity of positional ID modification, while Selective Routing adaptively routes between invariant and sensitive LMs to handle real-world scenarios.
+Together, these methods demonstrate improved performance and reliability on LitM, KGQA, and MMLU benchmarks.
 7 Limitations
@@
 Appendix A Appendix
-A.1 Exact values of the lost-in-the middle benchmark
+A.1 Full results on the Lost in the Middle Benchmark
 (a) Results with index bias (indexed by numbers). (Example input at Appendix Fig. 7)
@@
 1) due to its low performance.
+Total ndoc (segments) 10 20 30
+Gold idx at: 0 4 9 avg. 0 4 9 14 19 avg. 0 4 9 14 19 24 29 avg.
+Indexing bias present
+Original 52.8 51.1 46.7 50.2 55.0 50.7 50.5 50.2 47.5 50.7 56.1 50.8 51.5 51.1 50.8 50.5 48.9 51.4
+PCW 12.0 11.9 12.1 12.0 03.4 03.7 03.8 03.9 03.6 05.1 02.1 02.1 02.0 01.9 02.0 02.2 02.0 02.0
+Set-Based Prompting 40.8 40.7 40.8 40.8 25.6 25.8 25.7 25.5 25.3 25.6 15.8 15.9 16.1 16.0 16.1 15.7 15.8 15.9
+PINE 59.2 56.8 57.7 57.9 56.2 55.1 54.5 54.2 55.3 55.5 56.1 53.6 53.3 53.9 53.2 53.7 54.5 54.0
+RoToR-lexical 63.6 60.5 59.1 61.1 52.6 58.6 57.8 56.8 58.6 57.6 64.6 58.9 56.9 56.2 57.1 56.2 57.1 58.1
+RoToR-reversed lexical 61.5 60.8 60.6 61.0 60.8 58.6 59.7 60.5 59.3 60.0 61.1 57.5 58.2 58.6 58.3 59.3 59.2 58.9
+RoToR-reranking 61.4 61.4 61.7 61.5 62.3 59.2 59.1 59.4 59.8 60.2 62.2 58.6 58.4 58.6 58.3 58.5 59.8 59.2
+RoToR-freq 62.8 61.1 59.5 61.1 62.9 58.8 56.7 57.4 58.0 59.1 61.7 58.2 56.9 56.1 56.4 55.4 56.8 57.4
+Indexing bias removed (main paper)
+Original 54.7 53.0 50.2 52.6 54.8 52.6 52.8 52.4 51.0 52.7 55.6 51.5 52.4 52.8 52.1 52.3 53.0 52.8
+PCW 12.4 11.9 12.2 12.2 03.7 04.0 04.0 04.0 03.9 03.9 02.3 01.8 02.0 02.0 02.1 02.0 02.0 02.0
+Set-Based Prompting 42.5 42.5 42.5 42.5 26.3 26.3 26.3 26.3 26.3 26.3 14.1 14.1 14.1 14.1 14.1 14.1 14.1 14.1
+PINE 58.6 58.8 59.0 58.8 56.2 55.7 55.5 55.7 55.5 55.7 54.2 54.8 54.3 53.7 54.8 54.2 54.0 54.3
+RoToR-lexical 61.4 61.6 61.6 61.5 61.4 59.8 59.6 59.6 59.8 60.0 59.2 59.5 59.4 59.1 59.0 59.3 59.1 59.2
+RoToR-reversed lexical 61.6 61.8 61.8 61.8 58.9 59.3 58.8 58.6 58.7 58.8 57.9 58.2 57.9 57.4 57.9 57.6 57.5 57.8
+RoToR-reranking 61.2 61.4 61.2 61.3 60.9 61.0 61.2 61.2 61.2 61.1 60.9 60.7 60.7 60.7 60.8 60.8 60.7 60.8
+RoToR-freq 61.0 61.1 61.1 61.1 60.4 60.3 58.6 60.2 60.0 59.9 59.3 60.4 59.7 59.5 59.5 59.6 59.2 59.6
+[…]
 ndoc = 10 20
-gold idx = 0 4 9 avg. 0
-Orig. 66.2% 65.7% 65.7% 65.9% 65.2%
-PINE 67.9% 67.8% 67.5% 67.7% 65.9%
-RoToR-lex. 69.6% 69.5% 69.3% 69.5% 67.6%
-Table 5: Reporting the performance with Llama-3.1-70B-Instruct on a subset of lost-in-the middle benchmark (without index bias) performance.
+gold idx = 0 4 9 avg. 0 4 9 14 19 avg.
+Orig. 66.2 65.7 65.7 65.9 65.2 64.3 65.0 66.2 64.8 65.2
+PINE 67.9 67.8 67.5 67.7 65.9 65.7 65.9 65.8 65.5 66.1
+RoToR-reranking 68.9 69.0 68.8 68.9 67.5 67.5 67.7 67.5 67.6 67.8
+RoToR-lexical 69.6 69.5 69.3 69.5 67.6 67.8 67.8 67.7 67.9 68.0
+Table 6: Reporting the performance with the Llama-3.1-70B-Instruct model on a subset of lost-in-the middle benchmark (without index bias) performance.
 We see that the gains are consistant with the models we mainly investigated (Llama-3.1-8B-Instruct)
-Tab. 1 in the main paper reports the full number to the lost in the middle benchmark, and Fig. 5 visualizes them in graph format. Table  5 shows results in different model (Llama-3.1-70B-Instruct).
+Impact of removing index bias on LitM
+Tab.5 presents the full results on the Lost in the Middle (LitM) benchmark, comparing scenarios where indexing bias is present versus removed. Fig.5 provides a visual representation of these results.
+As shown in Appendix Tab.5, invariant LMs exhibit stable performance regardless of the gold index, especially when index bias is removed (as described in Sec.4.2; see also Appendix Fig.5(b)).
+However, when index bias is present, performance fluctuations are observed (Appendix Fig.5(a)).
+Notably, RoToR achieves the highest performance across all setups, demonstrating its effectiveness in mitigating positional bias in a zero-shot setting while maintaining overall performance.
+These findings suggest that index bias acts as an implicit source of additional positional bias and that invariant LMs benefit significantly from its removal.
+Results on other models.
+Table  6 shows results in different model (Llama-3.1-70B-Instruct). We can see that the trend continues for the 70B model, implying the robustness of RoToR.
 A.2 Illustration of the global sorting method
@@
 All inferences were done with a single NVIDIA RTX A6000 48GB GPU.
-For reproducibility, we fix the seed and use greedy decoding without any sampling for all experiments.
+Note that all of the baseline models including our method can be applied directly in a zero-shot, training-free manner.
+For reproducibility, we fix the seed and disabled random sampling (i.e., used greedy decoding), set the maximum number of new generated tokens to 500, and set the pad_token_id to the same value as the eos_token_id for all experiments.
 For all datasets we tested, we separate the input to 3 parts: prefix, parallel contexts, and suffix and feed them accordingly to the positional-invariant model.
@@
 A.3.2 Knowledge Graph Question Answering
-For evaluation with Mintaka and WebQSP, we follow the same setup as  Baek et al.
+For evaluation with Mintaka, we follow the same setup as  Baek et al.
 (2023).
@@
 We report the full variation of Selective Routing results on the investigated $\alpha$ value at Tab.
-6.
+7.
 A.6 Replication details on the runtime experiment
@@
 The runtime of each experiment was measured on an ASUS ESC8000-E11 server featuring dual 4th Gen Intel Xeon Scalable processors, 64 CPU threads, 1.1 TB of RAM across 32 DIMM slots, and 8 NVIDIA A6000 GPUs with 48 GB of memory each.
-Except for the experiments on Llama-3.1-70B-Instruct, we only use a single A6000 GPU for all of the experiments.
+We Except for the experiments on Llama-3.1-70B-Instruct, we only use a single A6000 GPU for all of the experiments.
 A.7 Input data examples
-To get a better understanding of the input and output format of the lost-in-the middle dataset and structured knowledge grounding dataset, we provide example inputs for each data format.
+To illustrate the input and output formats used in our experiments, we provide example inputs for the Lost-in-the-Middle (LitM), Knowledge Graph Question Answering (KGQA), and MMLU datasets.
+For experiments using the Qwen-Chat model, special tokens were adjusted accordingly.
+While the example prompts are based on the Llama-3.1-8B-Instruct model, the specific differences in token usage for the Qwen-Chat variants can be observed by comparing the prompts in Fig.
+8 and Fig.
+9.
+This adjustment is consistently applied across all datasets.
+Note that no special tokens are added for the MMLU benchmark, which aligns with the lm-evaluation harness setup.
 lost in the middle Prefix:
@@
 <|end_header_id|>
-Figure 8: Example input for the lost in the middle dataset, without indexing by numbers.
-WebQSP Prefix:
-<|begin_of_text|><|start_header_id|>system<|end_header_id|>
-Below are the facts in the form of the triple meaningful to answer the question.
-Answer the given question in a JSON format, such as "Answer": "xxx".
-Only output the JSON, do NOT say any word or explain.
-<|eot_id|><|start_header_id|>user<|end_header_id|> Parallel texts:
-(Ferdinand Magellan, occupation, seafarer)
-(Ferdinand Magellan, country of citizenship, Kingdom of Portugal) (Ferdinand Magellan, date of birth, time: +1480-00-00) (Ferdinand Magellan, instance of, human) (Ferdinand Magellan, date of death, time: +1521-04-27) (Ferdinand Magellan, sex or gender, male) (Ferdinand Magellan, killed by, Lapulapu) Suffix:
-Question: what country did magellan represent?, Answer: <|eot_id|><|start_header_id|>assistant <|end_header_id|>
-Gold Answer(s):
-(‘Reino de Portugal’, ‘Portuguese Kingdom’, ‘Kingdom of Portugal’)
-Example generated output:
-{"Answer": "Portugal"} (Parsed to: Portugal)
-Figure 9: Example input for the WebQSP dataset.
+Figure 8: Example input for the lost in the middle dataset, without indexing by numbers. Prompt for the Llama-3.1-8B-Instruct model.
+lost in the middle no indexing (Qwen variant) Prefix:
+<|im_start|>system
+You are a helpful, respectful and honest assistant.
+Always answer as helpfully as possible, while being safe.
+Please ensure that your responses are socially unbiased and positive in nature.
+If a question does not make any sense, or is not factually coherent, explain why instead of answering something not correct.
+If you don’t know the answer to a question, please don’t share false information.<|im_end|><|im_start|>user Write a high-quality answer for the given question using only the provided search results (some of which might be irrelevant).
+Parallel texts:
+[Document Title: Thorax] when deep breaths are attempted. Different people …
+[Document Title: Chest pain] present with chest pain, and carry a significantly higher … Suffix:
+Question: for complaints of sudden chest pain patients should take a<|im_end|>
+<|im_start|>assistant
+Figure 9: Example input for the lost in the middle dataset, without indexing by numbers, prompt for the Qwen1.5-Chat model.
 Mintaka Prefix:
@@
 Figure 10: Example input for the Mintaka dataset.
+MMLU Prefix:
+The following are multiple choice questions (with answers) about moral disputes.
+Norcross agrees that if a being is incapable of moral reasoning, at even the most basic level, then it cannot be Parallel texts:
+A. a being of value.
+B. an object of moral sympathy. C. a moral agent. D. a moral patient. Suffix:
+Answer:
+Figure 11: Example input for the MMLU benchmark.
 Appendix B Why is removing index bias an important detail for invariant models to be effective?
@@
 Appendix C Statisticial significance before and after shuffling segments
-We have calculated the paired one-tailed t-tests performed on both the baseline (“original”) model and our proposed method (RoToR) at Tab.
-8 using the results at Tab.
-7 on whether the performance differences between initial ordering v.s.
-shuffled ordering are statistically significant on the original model and ours.
-Note that the performance on the Lost-in-the Middle (LitM) was not reported since the LitM benchmark didn’t provide initial ordering information.
-Specifically, we test whether the mean performance difference, (Before Shuffle - After Shuffle), deviates significantly from zero.
-As shown in Table 8, the results indicate that while the original model does exhibit a statistically significant drop in performance upon shuffling, our RoToR method does not, suggesting increased robustness to segment-order perturbations.
-[^8] The calculation were validated using the online t-test calculator.
-https://www.mathportal.org/calculators/statistics-calculator/t-test-calculator.php
+We conducted paired two-tailed t-tests (Table 9) for both the baseline (“original”) model and our proposed method (RoToR), using the results in Table 8.
+Our goal was to determine whether the performance differences between the initial ordering and shuffled ordering are statistically significant.
+We excluded the Lost-in-the-Middle (LitM) dataset because it does not provide an initial ordering.
+Specifically, the tests evaluate whether the mean performance difference (Before Shuffle - After Shuffle) significantly deviates from zero.
+For KGQA, we selected the F1 score as the representative metric among the three available, gathering data points from various task configurations and different models.
+For MMLU, the results are based on our RoToR variant with selective routing.
+As shown in Table 9, the original model shows a statistically significant drop in performance when the segments are shuffled, while RoToR does not, indicating increased robustness to segment-order perturbations.
+[^8] All statistical calculations were validated using an online t-test calculator: https://www.mathportal.org/calculators/statistics-calculator/t-test-calculator.php
 Original Model RoToR-lexical
-Before Shuff. After Shuff. Diff. Before Shuff. After Shuff. Diff.
-MMLU 68.3 65.5 2.8 68.5 65.7 2.8
-WebQSP, ndoc 30 57.6 57.0 0.6 58.8 58.7 0.1
-WebQSP, ndoc 50 57.5 56.5 1.0 58.8 58.7 0.1
-Mintaka, ndoc 30 48.7 47.9 0.8 51.2 50.9 0.3
-Mintaka, ndoc 50 48.6 48.1 0.5 50.8 50.7 0.1
+Before
+Shuff.
+After
+Shuff.
+Diff.
+Before
+Shuff.
+After
+Shuff.
+Diff.
+Mintaka, Llama3.1-8B-Instruct, ndoc=30 51.9 51.0 0.9 54.1 53.8 0.3
+Mintaka, Llama3.1-8B-Instruct, ndoc=50 51.7 51.0 0.7 53.6 53.5 0.1
+Mintaka, Qwen1.5-4B-Chat, ndoc=30 34.9 34.7 0.2 35.7 35.5 0.2
+Mintaka, Qwen1.5-4B-Chat, ndoc=50 35.8 35.0 0.8 36.2 36.1 0.1
+Mintaka, Qwen1.5-7B-Chat, ndoc=30 35.4 35.0 0.4 37.7 37.7 0
+Mintaka, Qwen1.5-7B-Chat, ndoc=50 35.7 35.5 0.2 38.0 38.0 0
+MMLU, Llama3.1-8B-Instruct 68.3 65.5 2.8 68.5 65.7 2.8
+MMLU, Qwen1.5-4B-Chat 53.6 52.6 1 53.7 52.6 1.1
+MMLU, Qwen1.5-7B-Chat 60.1 58.6 1.5 60.1 58.8 1.3
 Table 8: Performance of the Original model and RoToR before and after shuffling.
 Original RoToR
-Degrees of Freedom 4
-Mean Difference 1.14 0.68
-t-Statistic 3.01 1.43
-Critical Value 2.132
+Degrees of Freedom 8
+Mean Difference 0.94 0.66
+t-Statistic 3.71 2.23
+Critical Value 2.306
 Statistically Significant Yes No
-Table 8: Paired one-tailed t-test results comparing the original model and ours.
+Table 9: Paired two-tailed t-test results comparing the original model and ours.
 Derivation for the original model.
-Let the five paired differences (Before -- After) be $\{d_{1},d_{2},d_{3},d_{4},d_{5}\}$.
-Mean Difference: $\bar{d}=\tfrac{1}{5}\sum_{i=1}^{5}d_{i}$.
-In this case, $\bar{d}\approx 1.14\%.$ Sample Standard Deviation: $s_{d}=\sqrt{\frac{1}{n-1}\sum_{i=1}^{n}(d_{i}-\bar{d})^{2}}\approx 0.8476.$ Standard Error (SE): $\mathrm{SE}=\tfrac{s_{d}}{\sqrt{n}}\approx 0.379.$ t-Statistic: $t=\tfrac{\bar{d}}{\mathrm{SE}}\approx 3.01,\ (df=4).$ Since the critical value at $df=4$ and $\alpha=0.05$ is $2.132,$ we have $3.01>2.132$ Therefore, the difference is statistically significant.
+Let the nine paired differences (Before -- After) be $\{d_{1},d_{2},d_{3},...d_{8},d_{9}\}$.
+Mean Difference: $\bar{d}=\tfrac{1}{9}\sum_{i=1}^{9}d_{i}$.
+In this case, $\bar{d}\approx 0.9444\%.$ Sample Standard Deviation: $s_{d}=\sqrt{\frac{1}{n-1}\sum_{i=1}^{n}(d_{i}-\bar{d})^{2}}\approx 0.7632.$ Standard Error (SE): $\mathrm{SE}=\tfrac{s_{d}}{\sqrt{n}}\approx 0.2544.$ t-Statistic: $t=\tfrac{\bar{d}}{\mathrm{SE}}\approx 3.7124,\ (df=8).$ Since the critical value at $df=8$ and $\alpha=0.05$ is $2.306,$ we have $3.71>2.306$ Therefore, the difference is statistically significant.
 Derivation for the RoToR model.
-Under the same procedure, $\bar{d}\approx 0.68\%,\ s_{d}\approx 1.0628,\ \mathrm{SE}\approx 0.475.$ t-Statistic: $t\approx 1.4307.$ Since $1.4307<2.132,$ there is no significant difference in performance before and after shuffling for RoToR.
+Under the same procedure, $\bar{d}\approx 0.6556\%,\ s_{d}\approx 0.8833,\ \mathrm{SE}\approx 0.2944.$ t-Statistic: $t\approx 2.2265.$ Since $2.2265<2.306,$ there is no significant difference in performance before and after shuffling for RoToR.
 Conclusion.

### v3

@@
 While zero-shot order-invariant LMs have been proposed to solve this issue, their success on practical listwise problems has been limited.
-In this work, as a first contribution, we identify and overcome two limitations to make zero-shot invariant LMs more practical: (1) training and inference distribution mismatch arising from modifying positional ID assignments to enforce invariance, and (2) failure to adapt to a mixture of order-invariant and sensitive inputs in practical listwise problems.
+In this work, as a first contribution, we identify and overcome two limitations to make zero-shot invariant LMs more practical: (1) training and inference distribution mismatch arising from modifying positional ID assignments to enforce invariance, and (2) failure to adapt to mixture of order-invariant and sensitive inputs in practical listwise problems.
 Then, to overcome these issues we propose (1) RoToR, a zero-shot invariant LM for genuinely order-invariant inputs with minimal modifications of positional IDs, and (2) Selective Routing, an adaptive framework that handles both order-invariant and order-sensitive inputs in listwise tasks.
@@
 Soyoung Yoon1thanks:   Work done during an internship at Channel Corporation.
-Dongha Ahn1    Youngwon Lee1    Minkyu Jung2    HyungJoo Jang2    Seung-won Hwang1thanks:   Corresponding author.
+Dongha Ahn12    Youngwon Lee1    Minkyu Jung2    HyungJoo Jang2    Seung-won Hwang1thanks:   Corresponding author.
 1Seoul National University 2Channel Corporation {soyoung.yoon, seungwonh}@snu.ac.kr
@@
 (2024); Zheng et al.
-(2024a) also investigate the importance of positional bias, extending the domain to structured knowledge grounding (SKG) tasks Zhao et al.
+(2024a) also investigate the impact of positional bias, extending the domain to structured knowledge grounding (SKG) tasks Zhao et al.
 (2023); Zhou et al.
@@
 (2024); Cai et al.
-(2023) models, they leave decoder-only models, which account for most of the current LLMs, for more consideration.
+(2023) models, they leave decoder-only models, which account for the current frontier LLMs, for more consideration.
 In contrast to transformer encoders that use bidirectional attention which is invariant by nature Lee et al.
@@
 (2023); Zhu et al.
-(2023), we focus on zero-shot approaches for practicality, namely PCW, Set-Based Prompting Ratner et al.
-(2023); McIlroy-Young et al.
+(2023), we focus on zero-shot approaches for practicality, namely PCW Ratner et al.
+(2023), Set-Based Prompting McIlroy-Young et al.
 (2024), and PINE Wang et al.
@@
 3 Methodology
-Figure 2: For example input ‘‘Given Apple Banana Orange which one is most related to the color red?
-A: It is ...’’, demonstration of how PINE and RoToR (ours) modify the attention mask and positional IDs for each query-key combination, resulting in segment-wise order invariance.
-Assume each block represents a single token, and blocks are colored in positional ID assignment.
-White (no-block) area indicate masked attention.
+Figure 2: Attention mask and positional ID modifications for segment-wise order invariance using example input ‘‘Given Apple Banana Orange...’’.
+Each block represents a single token arranged by positional ID assignment.
+In practice, tokens retain original positions but receive reassigned positional IDs as shown.
+White areas indicate masked attention.
+Key difference: PINE requires per-query sorting while RoToR reuses one global ordering across all queries.
 3.1 Baseline: Order-invariant causal LMs
@@
 Bidirectional processing with Q-K similarity
-A more recent work, PINE (Wang et al., 2024) has addressed these issues through a bidirectional attention mechanism, by letting each segment attend to all other segments.
-However, to allow this within decoder-only models with causal attention and still achieve order invariance, PINE modifies the multi-head self-attention procedure to create an ‘illusion’: it treats each query segment as if it were the final segment in the input list, so that it can attend to all other segments as keys.
-For each query segment, the ordering among the key segments is determined by their attention scores (without positional embeddings, $\mathrm{Attn}_{\mathrm{NoPoS}}$), ensuring that segments with stronger relevance to the query appear closer.
-Meanwhile, the tokens within a query segment still follow causal ordering.
-Fig.
-2 illustrates this with the input [T1‘‘Given’’, S1[‘‘Apple’’], S2[‘‘Ban’’, ‘‘ana’’], S3[‘‘Orange’’], T6‘‘which one’’, ..
-T10‘‘red?’’].
-Prefix tokens ‘‘Given’’ and suffix tokens ‘‘which one ..
-red?’’ remain in their original positions and follow normal causal attention, equally attending to all segments.
-In contrast, the segments in the middle (S1 - S3) require the order-invariant mechanism.
-PINE dynamically re-assigns positional IDs depending on a token’s role as a query or a key, instead of using a single fixed positional ID.
-For Case 1: When “ana” (T4, S2) acts as a query, S2 is assigned the largest position IDs (5) so it can attend to all previous segments.
-Within S2, the tokens “ban” (4) and “ana” (5) still maintain their local causal order (i.e., smaller to larger positions).
-Case 2: When “ana” serves as a key, the placement of S2 depends on its $\mathrm{Attn}_{\mathrm{NoPoS}}$ scores with the query segment.
-If “banana” (S2) is more relevant to the query segment “apple” (S1) than “orange” (S3), (i.e., $\mathbf{Attn_{NoPos}}$(S1, S2) > $\mathbf{Attn_{NoPos}}$ (S1, S3)) S2 is placed nearer to S1, with “ban” and “ana” preserving their internal sequence.
-For prefix, suffix, and generated tokens, they do not participate in order invariance, so they are always placed at their standard causal positions.
-However, the same dynamic reordering applies to segment tokens (T2–T5) when computing attention scores of suffix (T6–T10) and the generated token (T11) as query.
+A more recent work, PINE (Wang et al., 2024) has addressed these issues through a bidirectional attention mechanism that allows each segment to attend to all other segments.
+To achieve this within decoder-only models while maintaining order invariance, PINE dynamically modifies positional IDs based on whether a token acts as a query or key in the attention computation.
+The key insight is that PINE creates an “illusion” for each query segment: it assigns the query segment the largest positional IDs among all segments, enabling it to attend to all other segments bidirectionally.
+The ordering of key segments is then determined by their relevance scores computed without positional embeddings ($\mathrm{Attn}_{\mathrm{NoPoS}}$), ensuring that more relevant segments appear closer to the query.
+Consider the example in Fig.
+2 with input [T1‘Given’’, S1[Apple’’], S2[Ban’’, ana’’], S3[Orange’’], T6‘which one’’, ..
+T10‘red?’’].
+The prefix token ‘Given’’ (T1) and suffix tokens ‘which one ..
+red?’’ (T6-T10) maintain their original positions and follow standard causal attention.
+For the segments S1-S3, PINE applies its order-invariant mechanism:
+[…]
+This dynamic reassignment occurs for every attention computation: each query token sees a different positional arrangement of the key tokens, determined by their relevance scores.
+Prefix, suffix, and generated tokens do not participate in this reordering and always maintain their standard causal positions.
+However, when these non-segment tokens act as queries, they still see the segments reordered by their relevance scores.
 Figure 3: Comparing the ordering of 5 segments (S1 - S5) of PINE (Wang et al., 2024; left) and RoToR (Ours; right).
@@
 3.2 RoToR: minimal OOD from positional ID assignments
+Challenges with PINE
 While PINE achieves order-invariance by contextualization across segments, its query-specific ordering scheme introduces (1) significant train-test behavior discrepancy as well as (2) unnecessary complexity and numerical instability, which limits its scalability.
@@
 bfloat16 numeric type lacks precision to distinguish these values, leading to non-determinism originating from several tied values.
-The outcome may depend on the initial ordering; to address these problems stemming from query-dependent ordering, we instead propose RoToR (Fig.
-3), which uses one global ordering that is not a function of the initial ordering of segments (e.g., canonical ordering by lexical sorting) and assign IDs for tokens in different segments based on circular arrangement.
+The outcome may then depend on the initial ordering.
+Motivation & Theoretical Foundation
+While investigating ways to overcome the limitations of PINE, our central goal is to preserve order invariance while minimizing the complexity of the re-assignment of positions.
+We reason that defining a single global ordering scheme, not necessarily relying on attention scores, and re-using them across all queries can solve the problems stemming from query-dependent ordering.
+A circular assignment of a global order seems as a practical solution.
+The idea of using global sorting to achieve ordering invariance has been studied in set/graph ML domain (Murphy et al., 2019a, b), but to the best of our knowledge, using circular assignment of the global ordering, and the application to pre-trained language models, are our novel contributions.
+As a result, we propose RoToR(Fig.
+3), which uses one global ordering that is not a function of the initial ordering of segments (e.g., canonical ordering by lexical sorting) and assigns IDs for tokens in different segments based on circular arrangement.
 Global ordering
@@
 For example, in Figure 3, due to the global ordering, segments $S_{5}$ and $S_{2}$ are always placed in adjacent positions with RoToR (right side), while it is not satisfied and constantly changed with PINE (left side).
-We consider three separate global sorting algorithm to be used in RoToR: (1) simple lexicographical sorting which can be obtained with minimal overhead based on tokenized sequence of segments, (2) using a pointwise reranker333castorini/monot5-base-msmarco-10k to score relevancy of each row with respect to the question, or (3) simple frequency-based sorting which normalizes token ids based on the inverse frequency of each token (Details at Fig.
+We consider three separate global sorting algorithm to be used in RoToR: (1) simple lexicographical sorting which can be obtained with minimal overhead based on tokenized sequence of segments, (2) using a pointwise reranker Nogueira et al.
+(2020)333castorini/monot5-base-msmarco-10k to score relevancy of each row with respect to the question, or (3) simple frequency-based sorting which normalizes token ids based on the inverse frequency of each token (Details at Appendix Fig.
 6).
-Empirically, we find that using simple lexicographical sorting is sufficient for bringing improvements over PINE.
+Empirically, we find that using simple lexicographical sorting is sufficient to obtain improvements over PINE.
 Circular arrangement
@@
 Computational overhead
-By adopting global sorting, we can avoid extra attention scores computation and thus improve efficiency.
-While the cost with PINE to obtain hidden states is $\mathcal{O}(n^{2}d+nk\log{k})$ for input length $n$, hidden dimension $d$, and $k$ segments (Wang et al., 2024), our method with lexicographical sorting achieves $\mathcal{O}(nk\log{k})$ cost, being more efficient and faster (lightweight) than PINE.
-We empirically validate that our method is much faster than PINE as $k$ increases, at Tab.
-4.
+We report only operations executed beyond vanilla self-attention cost $O(n^{2}d)$, where $n$ is total input length, $d$ is hidden dimension, and $k$ is the number of segments.
+PINE requires two additional operations: (1) computing attention scores without rotary position embeddings ($\mathcal{O}(n^{2}d)$) and (2) sorting $k$ segments for each query token ($\mathcal{O}(nk\log k)$), totaling $\bm{\mathcal{O}(n^{2}d+nk\log k)}$ (Wang et al., 2024) [^4] The PINE paper reports $\mathcal{O}(nk\log k)$ by absorbing the $\mathcal{O}(n^{2}d)$ term into baseline; we expose it explicitly for fair comparison..
+In contrast, our lexicographical sorting requires only a single global sort of $k$ segments ($\mathcal{O}(k\log k)$), each with length $\mathcal{O}(n)$, achieving $\bm{\mathcal{O}(nk\log k)}$ and eliminating the expensive $\mathcal{O}(n^{2}d)$ term entirely.
+This can be further optimized to $\bm{\mathcal{O}(nk)}$ using radix sort.
+[^5] https://en.wikipedia.org/wiki/Radix_sort We empirically validate significantly faster performance than PINE as $k$ increases (Tab.
+4).
 3.3 Selective Routing for handling order-sensitive inputs
-Figure 4: Illustration of Selective Routing (Sec. A.5).
+Figure 4: Illustration of Selective Routing (Sec. E).
 Since many practical benchmarks such as MMLU involves semi-invariant inputs, we propose a routing mechanism that uses the order-invariant model in conjunction with the standard causal model for further applicability.
@@
 $\alpha$ is a hyperparameter that controls how strongly the original model is favored, which was selected as 0.2 according to hyperparameter search on the validation subset (Appendix Sec.
-A.5).
+E).
 4 Experiment setup
@@
 RoToR-Freq. 61.0 61.1 61.1 60.4 60.3 58.6 60.2 60.0 59.3 60.4 59.7 59.5 59.5 59.6 59.2
+Llama 3.1-70B-Instruct
+Original 66.2 65.7 65.7 65.2 64.3 65.0 66.2 64.8
+PINE 67.9 67.8 67.5 65.9 65.7 65.9 65.8 65.5
+RoToR 69.6 69.5 69.3 67.6 67.8 67.8 67.7 67.9
+RoToR-MonoT5 68.9 69.0 68.8 67.5 67.5 67.7 67.5 67.6
 Qwen1.5-4B-Chat
@@
 […]
-Llama-3.1-8B-Instruct Qwen1.5-4B-Chat Qwen1.5-7B-Chat
-N = 30 N = 50 N = 30 N = 50 N = 30 N = 50
-Method Acc. EM F1 Acc. EM F1 Acc. EM F1 Acc. EM F1 Acc. EM F1 Acc. EM F1
+Llama-3.1-8B-Instr. Llama-3.1-70B-Instr. Qwen1.5-4B-Chat Qwen1.5-7B-Chat Qwen1.5-72B-Chat
+Method Acc. EM F1 Acc. EM F1 Acc. EM F1 Acc. EM F1 Acc. EM F1
+$N\!=\!30$
 Initial, no shuffling of segments
-Original 50.2 44.0 51.9 50.0 44.0 51.7 30.7 27.9 34.9 31.6 28.6 35.8 31.5 27.8 35.4 31.7 28.0 35.7
-PINE 51.5 45.0 52.6 51.6 45.1 52.6 31.6 28.7 35.6 31.6 28.8 35.3 32.3 28.8 36.4 32.0 28.5 35.9
-RoToR 53.1 46.5 54.1 52.9 46.0 53.6 32.0 29.0 35.7 32.7 29.6 36.2 34.3 29.8 37.7 34.3 30.1 38.0
-RoToR-MonoT5 51.6 45.0 52.5 52.4 45.4 52.8 32.3 29.1 36.2 32.3 29.3 35.9 32.9 28.4 36.3 32.9 28.9 36.6
-RoToR-Freq. 52.6 46.1 53.7 53.1 46.4 53.7 32.3 29.2 36.0 32.3 29.2 35.9 33.7 29.5 37.2 33.5 29.5 37.2
+Original 50.2 44.0 51.9 61.6 57.7 63.6 30.7 27.9 34.9 31.5 27.8 35.4 41.4 37.7 43.7
+PINE 51.5 45.0 52.6 63.1 58.7 64.8 31.6 28.7 35.6 32.3 28.8 36.4 46.7 42.9 49.0
+RoToR 53.1 46.5 54.1 63.6 59.1 65.2 32.0 29.0 35.7 34.3 29.8 37.7 47.5 43.2 49.2
+RoToR-MonoT5 51.6 45.0 52.5 32.3 29.1 36.2 32.9 28.4 36.3
+RoToR-Freq. 52.6 46.1 53.7 32.3 29.2 36.0 33.7 29.5 37.2
 After shuffling segments, averaged over 3 seeds
-Original 49.5 43.3 51.0 49.7 43.5 51.0 30.1 27.5 34.7 30.3 27.6 35.0 31.4 27.3 35.0 31.6 27.9 35.5
-$\hookrightarrow$ stdev. ($\pm$) 0.07 0.14 0.17 0.34 0.28 0.46 0.41 0.34 0.43 0.26 0.24 0.35 0.26 0.28 0.29 0.40 0.56 0.42
-PINE 51.8 45.2 52.8 51.8 45.3 52.7 31.5 28.7 35.6 31.5 28.7 35.3 32.3 28.8 35.7 31.7 28.2 35.7
-$\hookrightarrow$ stdev. ($\pm$) 0.05 0.07 0.16 0.15 0.16 0.19 0.20 0.18 0.13 0.17 0.20 0.21 0.17 0.20 0.13 0.18 0.16 0.14
-RoToR 52.8 46.2 53.8 52.7 45.9 53.5 31.8 28.8 35.5 32.5 29.6 36.1 34.2 29.9 37.7 34.2 30.1 38.0
-$\hookrightarrow$ stdev. ($\pm$) 0.05 0.05 0.02 0.05 0.09 0.04 0.05 0.02 0.09 0.11 0.06 0.09 0.09 0.07 0.06 0.06 0.05 0.04
-RoToR-MonoT5 51.6 45.0 52.6 52.2 45.2 52.8 32.4 29.2 36.3 32.3 29.4 35.9 33.0 28.8 36.5 32.8 28.8 36.5
-$\hookrightarrow$ stdev. ($\pm$) 0.12 0.06 0.10 0.16 0.18 0.18 0.04 0.02 0.13 0.16 0.13 0.07 0.12 0.09 0.07 0.16 0.09 0.07
-RoToR-Freq. 52.5 45.9 53.5 53.1 46.4 53.7 32.3 29.3 36.0 32.4 29.3 36.1 33.8 29.6 37.4 33.7 29.6 37.4
-$\hookrightarrow$ stdev. ($\pm$) 0.10 0.15 0.11 0.02 0.07 0.03 0.13 0.16 0.09 0.09 0.04 0.06 0.04 0.00 0.09 0.04 0.16 0.22
-Table 2: Results on the Mintaka (KGQA) dataset on different models, with standard deviation of the average scores with $\pm$. N refers to number of top-k segments per query.
+Original 49.5 43.3 51.0 62.1 57.8 64.0 30.1 27.5 34.7 31.4 27.3 35.0 41.0 37.6 43.6
+$\hookrightarrow$ stdev. ($\pm$) 0.07 / 0.14 / 0.17 0.37 / 0.40 / 0.27 0.41 / 0.34 / 0.43 0.26 / 0.28 / 0.29 0.75 / 0.40 / 0.33
+PINE 51.8 45.2 52.8 63.3 58.8 64.9 31.5 28.7 35.6 32.3 28.8 35.7 46.9 43.3 49.2
+$\hookrightarrow$ stdev. ($\pm$) 0.05 / 0.07 / 0.16 0.13 / 0.04 / 0.10 0.20 / 0.18 / 0.13 0.17 / 0.20 / 0.13 0.18 / 0.20 / 0.20
+RoToR 52.8 46.2 53.8 63.5 59.1 65.3 31.8 28.8 35.5 34.2 29.9 37.7 47.4 43.1 49.1
+$\hookrightarrow$ stdev. ($\pm$) 0.05 / 0.05 / 0.02 0.11 / 0.07 / 0.08 0.05 / 0.02 / 0.09 0.09 / 0.07 / 0.06 0.06 / 0.04 / 0.07
+RoToR-MonoT5 51.6 45.0 52.6 32.4 29.2 36.3 33.0 28.8 36.5
+$\hookrightarrow$ stdev. ($\pm$) 0.12 / 0.06 / 0.10 0.04 / 0.02 / 0.13 0.12 / 0.09 / 0.07
+RoToR-Freq. 52.5 45.9 53.5 32.3 29.3 36.0 33.8 29.6 37.4
+$\hookrightarrow$ stdev. ($\pm$) 0.10 / 0.15 / 0.11 0.13 / 0.16 / 0.09 0.04 / 0.00 / 0.09
+$N\!=\!50$
+Initial, no shuffling of segments
+Original 50.0 44.0 51.7 62.6 58.5 64.5 31.6 28.6 35.8 31.7 28.0 35.7 42.1 38.7 44.5
+PINE 51.6 45.1 52.6 64.1 59.8 65.8 31.6 28.8 35.3 32.0 28.5 35.9 48.0 44.1 49.9
+RoToR 52.9 46.0 53.6 64.6 60.0 66.2 32.7 29.6 36.2 34.3 30.1 38.0 48.4 44.3 50.3
+RoToR-MonoT5 52.4 45.4 52.8 32.3 29.3 35.9 32.9 28.9 36.6
+RoToR-Freq. 53.1 46.4 53.7 32.3 29.2 36.1 33.5 29.5 37.2
+After shuffling segments, averaged over 3 seeds
+Original 49.7 43.5 51.0 62.8 58.5 64.5 30.3 27.6 35.0 31.6 27.9 35.5 42.1 38.9 44.7
+$\hookrightarrow$ stdev. ($\pm$) 0.34 / 0.28 / 0.46 0.29 / 0.28 / 0.05 0.26 / 0.24 / 0.35 0.40 / 0.56 / 0.42 0.30 / 0.40 / 0.35
+PINE 51.8 45.3 52.7 64.3 59.8 65.9 31.5 28.7 35.3 31.7 28.2 35.7 48.0 44.3 50.0
+$\hookrightarrow$ stdev. ($\pm$) 0.15 / 0.16 / 0.19 0.16 / 0.15 / 0.14 0.17 / 0.20 / 0.21 0.18 / 0.16 / 0.14 0.02 / 0.04 / 0.05
+RoToR 52.7 45.9 53.5 64.5 60.0 66.1 32.5 29.6 36.1 34.2 30.1 38.0 48.3 44.3 50.3
+$\hookrightarrow$ stdev. ($\pm$) 0.05 / 0.09 / 0.04 0.02 / 0.02 / 0.01 0.11 / 0.06 / 0.09 0.06 / 0.05 / 0.04 0.05 / 0.09 / 0.05
+RoToR-MonoT5 52.2 45.2 52.8 32.3 29.4 35.9 32.8 28.8 36.5
+$\hookrightarrow$ stdev. ($\pm$) 0.16 / 0.18 / 0.18 0.16 / 0.13 / 0.07 0.16 / 0.09 / 0.07
+RoToR-Freq. 53.1 46.4 53.7 32.4 29.3 36.1 33.7 29.6 37.4
+$\hookrightarrow$ stdev. ($\pm$) 0.02 / 0.07 / 0.03 0.09 / 0.04 / 0.06 0.04 / 0.16 / 0.22
+Table 2: Mintaka (KGQA) results, with Initial and After-shuffle settings, across different model parameter size and backbones.
+$N$ refers to the number of top-k segments per query.
+Rows with $\hookrightarrow$ stdev.” report standard deviation over 3 seeds.
 Llama-3.1-8B-Instruct Qwen1.5-4B-Chat Qwen1.5-7B-Chat
@@
 $\hookrightarrow$ + S.R. 68.5 65.3 65.8 $\pm$ 0.8 53.7 52.3 52.6 $\pm$ 0.6 60.0 57.3 58.6 $\pm$ 0.8
-RoToR + S.R. (Optim.) 75.0 71.9 72.7 $\pm$ 1.0 61.8 60.1 61.1 $\pm$ 1.0 68.1 66.2 67.2 $\pm$ 0.7
+RoToR + S.R. (Oracle) 75.0 71.9 72.7 $\pm$ 1.0 61.8 60.1 61.1 $\pm$ 1.0 68.1 66.2 67.2 $\pm$ 0.7
 Table 3: Improving applicability to general listwise tasks (MMLU, N=4) with Selective Routing (S.R), which includes both order-invariant and order-sensitive examples.
-RoToR with Selective Routing shows improved performance and stability across re-orderings of input, and its high upper-bound (Optim.) implies the potential for further improvements.
+Init./Rev.
+refer to original/reversed orderings, Avg.
+is the average selection ratio across all (4!-1) re-orderings with standard deviation.
+S.R (Oracle) represents the upper bound with perfect routing accuracy.
+RoToR with Selective Routing shows improved performance and stability across input re-orderings.
 4.1 Baselines
@@
 (2024).
-We use the LLaMA 3.1 AI (2024) 8B-Instruct444meta-llama/Meta-Llama-3.1-8B-Instruct, Qwen1.5-4B-Chat and Qwen1.5-7B-Chat [^5] Qwen/Qwen1.5-4/7B-Chat as our backbone model for all of our experiments.
-As our method doesn’t need training, a single A6000 GPU was sufficient to run all of the experiments except for the Llama-3.1-70B-Instruct model (Appendix, Tab.
-6).
+We use the LLaMA 3.1 AI (2024) 8B-Instruct666meta-llama/Meta-Llama-3.1-8B-Instruct 70B-Instruct, Qwen1.5-4B-Chat, Qwen1.5-7B-Chat [^7] Qwen/Qwen1.5-4/7B-Chat, and Qwen1.5-72B-Chat as our backbone model for our experiments.
+As our method doesn’t need training, a single A6000 GPU was sufficient to run all of the experiments except for the Llama-3.1-70B-Instruct and Qwen1.5-72B-Chat model.
 We also conduct experiments on a subset of benchmarks (LitM and MMLU) on the runtime latency, perplexity, and collision rate of PINE and RoToR, to further validate our claims on Sec.
@@
 We experiment with three benchmarks involving real-world listwise input data.
-Examples of exact inputs and outputs are provided in Appendix A.7.
+Examples of exact inputs and outputs are provided in Appendix G.
 All reported scores are rounded to the nearest tenth, except for the standard deviation (rounded to the second decimal place).
@@
 Replication details and example dataset format are at Appendix Sec.
-A.3 and Fig.
+C and Fig.
 10.
@@
 Experiments on LitM found that eliminating the effect of index bias is another important detail for measuring true order robustness: (Appendix Sec.
-B).
+H).
 Thus, we report experiments with index bias eliminated.
@@
 8 and Sec.
-A.1.
+A.
 MMLU The Massive Multitask Language Understanding (MMLU) benchmark Hendrycks et al.
@@
 Effectiveness of RoToR
-We observe that shuffling input segments leads to non-trivial performance degradations in the original model, which exhibits a statistically significant performance drop on our experimented dataset (two-tailed t-test, p < 0.05, Appendix C).
+We observe that shuffling input segments leads to non-trivial performance degradations in the original model, which exhibits a statistically significant performance drop on our experimented dataset (two-tailed t-test, p < 0.05, Appendix I).
 In contrast, our proposed RoToR model does not show a statistically significant difference in performance before and after shuffling, indicating that it is more robust against such perturbations.
@@
 Improvements from PINE
-Runtime latency (s)
-Perple-
-xity ($\downarrow$)
-Collision
-Rate($\downarrow$)
-MMLU LitM LitM
-data count 14015 2655 2655
-# segments 4 10 30 20 30
-PINE 7,371 18,551 41,664 6.91 42.3%
-RoToR 6,608 14,264 23,569 6.65 0 (None)
-Reduced % 10.4% 23.1% 43.4%
-Table 4: Compared to PINE, RoToR with lexical sort is faster, has lower perplexity, and have zero collision rate between segments. We use a subset of LitM with index bias removed and gold index = 0.
+Model Benchmark PINE RoToR Reduction
+(a) Overhead FLOPs, relative to original model
+Llama-3.1- 8B-Instruct MMLU, $N=4$ 0.59$\times$ 0.55$\times$ 7.6%
+LitM, $N=10$ 7.07$\times$ 4.81$\times$ 31.9%
+LitM, $N=30$ 22.43$\times$ 15.05$\times$ 32.9%
+Llama-3.1- 70B-Instruct KGQA, $N=30$ 1.27$\times$ 0.94$\times$ 26.0%
+KGQA, $N=50$ 1.82$\times$ 1.29$\times$ 29.0%
+Qwen1.5- 72B-Chat KGQA, $N=30$ 0.45$\times$ 0.01$\times$ 98.0%
+KGQA, $N=50$ 0.58$\times$ 0.03$\times$ 94.8%
+(b) End-to-end latency (s)
+Llama-3.1- 70B-Instruct LitM, $N=10$ 57,352 44,219 22.9%
+LitM, $N=20$ 87,091 58,680 32.6%
+Llama-3.1- 8B-Instruct MMLU, $N=4$ 7,371 6,608 10.4%
+LitM, $N=10$ 18,551 14,264 23.1%
+LitM, $N=30$ 41,664 23,569 43.4%
+(c) Perplexity & Collision rate, (on LitM)
+Llama-3.1- 8B-Instruct Perplexity ($N=20$) 6.91 6.65
+Collision rate ($N=30$) 42.3% 0 (None)
+Table 4: Unified efficiency comparison of RoToR vs.
+PINE, reporting (a) Additional FLOPs, (b) Latency, and (c) Perplexity & Collision rate.
+Columns list each metric for PINE and RoToR, and the relative reduction.
+Yellow rows separate sub-sections.
 Experiments against comparing RoToR with PINE (Tab.
-4) analyze the following: Runtime, scalability: Actual inference times (Appendix Sec.
-A.6) find that RoToR outperforms PINE substantially, with efficiency gains increasing alongside $n$.
+4) we analyze FLOPs:  [^8] We used the FlopsProfiler of the DeepSpeed library to measure FLOPs.
+RoToR consistently reduces the floating point operations overhead across segment counts and different model backbones compared to PINE.
+This is because RoToR does not require computing additional attention scores: it only performs tensor operations for circular arrangement.
+In contrast, PINE requires more cost due to attention-based reassignment.
+Runtime, scalability: Actual inference times (Appendix Sec.
+F) find that RoToR outperforms PINE substantially, with efficiency gains increasing alongside $n$.
 For instance, on LitM (30 docs), RoToR achieves a 43% reduction in total runtime.
@@
 Selective Routing improves the generalizability and extends the applicability on practical listwise tasks by adaptively handling order-sensitive inputs.
-The RoToR + Selective Routing (Optimal) performance on Tab.
+The RoToR + Selective Routing (Oracle) performance on Tab.
 3 was evaluated using a relaxed accuracy metric based on the union of predictions from the original and the invariant (RoToR-lexical) model.
 This improves significantly, which highlights the potential of Selective Routing for further accuracy gains through optimizing design choices on routing methods, which we plan to explore in future work.
+Additional analysis on the selection ratio of Selective Routing is reported at Appendix Section L.
 Impact of global ordering algorithm
@@
 1 show that any type of global sorting, with the use of circular assignment is superior than PINE, which relies on pairwise attention arrangements.
+Extension to LLMs, other scenarios
+Experiments on Llama-3.1-70B-Instruct and Qwen-1.5-72B-Chat for LitM and KGQA show consistent and significant improvements over both the original implementation and the PINE baseline, demonstrating RoToR’s generalizability to larger-scale LLMs.
+We further evaluate robustness on longer-context inputs (LongBench-2WikiMultiHopQA Bai et al.
+(2024) in Appendix Section J) and different task templates (KGQA template swap in Appendix Section K).
+Results indicate that our method retains benefits across longer input scenarios and diverse task templates, including those without explicit input format requirements.
 6 Conclusion
 Our work addresses order-invariance in listwise inputs by identifying core issues in distribution mismatch and adaptive handling of mixed inputs.
-Our proposed RoToR provides a stable zero-shot order-invariant solution that reduces the complexity of positional ID modification, while Selective Routing adaptively routes between invariant and sensitive LMs to handle real-world scenarios.
+Our proposed RoToR by modifying self-attention by global sorting and circular arrangement provides a stable zero-shot order-invariant solution that reduces the complexity of positional ID modification, while Selective Routing adaptively routes between invariant and sensitive LMs to handle real-world scenarios.
 Together, these methods demonstrate improved performance and reliability on LitM, KGQA, and MMLU benchmarks.
@@
 It would be beneficial to support more complex structures, such as ability to process multiple order-invariant contexts interleaved with serial text.
-Appendix A Appendix
-A.1 Full results on the Lost in the Middle Benchmark
+8 Acknowledgments
+We thank the Channel Corporation for providing GPU resources to run the experiments, and their AI team for providing helpful feedback.
+We thank the members of LDILab and Jinwoo Kim for their constructive comments and the anonymous reviewers for their valuable suggestions.
+We are also grateful to our former intern, Yezun Chung (KAIST CS), for assisting with experiments on early versions of the Selective Routing approach.
+This work was partly supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (No.
+RS-2021-II212068, Artificial Intelligence Innovation Hub) and Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) [NO.RS-2021-II211343, Artificial Intelligence Graduate School Program (Seoul National University)]
+Appendix
+Appendix A Full results on the Lost in the Middle Benchmark
 (a) Results with index bias (indexed by numbers). (Example input at Appendix Fig. 7)
@@
 […]
-ndoc = 10 20
-gold idx = 0 4 9 avg. 0 4 9 14 19 avg.
-Orig. 66.2 65.7 65.7 65.9 65.2 64.3 65.0 66.2 64.8 65.2
-PINE 67.9 67.8 67.5 67.7 65.9 65.7 65.9 65.8 65.5 66.1
-RoToR-reranking 68.9 69.0 68.8 68.9 67.5 67.5 67.7 67.5 67.6 67.8
-RoToR-lexical 69.6 69.5 69.3 69.5 67.6 67.8 67.8 67.7 67.9 68.0
-Table 6: Reporting the performance with the Llama-3.1-70B-Instruct model on a subset of lost-in-the middle benchmark (without index bias) performance.
-We see that the gains are consistant with the models we mainly investigated (Llama-3.1-8B-Instruct)
 Impact of removing index bias on LitM
@@
 These findings suggest that index bias acts as an implicit source of additional positional bias and that invariant LMs benefit significantly from its removal.
-Results on other models.
-Table  6 shows results in different model (Llama-3.1-70B-Instruct). We can see that the trend continues for the 70B model, implying the robustness of RoToR.
-A.2 Illustration of the global sorting method
+Appendix B Illustration of the global sorting method
 Figure 6: Illustration of ordering 7 rows by 2 different global sort options, (1) lexical sort based on token ids, or (2) reranking sort based on a reranker model (MonoT5-base in this case).
 (3) frequency based sorting applies lexical sorting, but the definitive token ids are mapped into the inverse frequency.
-We show the two different global sorting algorithms presented in our paper at Fig. 6.
-A.3 Details about preprocessing and evaluation of datasets
-A.3.1 General.
+We show the three different global sorting algorithms presented in our paper at Fig. 6.
+Appendix C Details about preprocessing and evaluation of datasets
+C.1 General.
 All inferences were done with a single NVIDIA RTX A6000 48GB GPU.
@@
 For PCW, we use the pcw_generate function, and we additionally utilized the RestrictiveTokensLogitsProcessor provided at the official PCW repository for MMLU classification, to have a similar setup with the log_likelihood option used for other models.
-A.3.2 Knowledge Graph Question Answering
+C.2 Knowledge Graph Question Answering
 For evaluation with Mintaka, we follow the same setup as  Baek et al.
@@
 For the row shuffling setup, we fix the seed to 0, 1, 2 on shuffling rows and report the average scores.
-A.3.3 Lost in the Middle
-Specifically, we use the dataset provided in the official repository666github.com/nelson-liu/lost-in-the-middle, use the same prompt as the llama 2 chat model with only the instruction tokens adjusted to llama 3 (removed [Inst] and changed to <|begin_of_text|> and etc.,), and evaluate using the best_supspan_em metric.
-A.3.4 MMLU
+C.3 Lost in the Middle
+Specifically, we use the dataset provided in the official repository999github.com/nelson-liu/lost-in-the-middle, use the same prompt as the llama 2 chat model with only the instruction tokens adjusted to llama 3 (removed [Inst] and changed to <|begin_of_text|> and etc.,), and evaluate using the best_supspan_em metric.
+C.4 MMLU
 We follow the publicly acknowledged lm-evaluation harness Gao et al.
@@
 We measure accuracy between the gold answer and the token with the highest likelihood (probability) among possible answer tokens [‘ A’, ‘ B’, ‘ C’, ‘ D’].
-A.4 Further impact scenarios on general conversation.
+Appendix D Further impact scenarios on general conversation.
 We shortly discuss about how this method may be applied to general conversational scenarios of LLMs.
@@
 Here, RoToR achieves a significant 7-9% average accuracy gain over the original LLM, very consistently across all setups (doc indexing and ndoc) for all choices of the ordering algorithm, with lower standard deviation than the original model.
-A.5 Selection of $\alpha$ for Selective Routing on MMLU
+Appendix E Selection of $\alpha$ for Selective Routing on MMLU
 $\alpha$ = no Selective Routing -0.5 -0.4 -0.3 -0.2 -0.1 0 0.1 0.2 0.3 0.4 0.5
@@
 We report the full variation of Selective Routing results on the investigated $\alpha$ value at Tab.
-7.
-A.6 Replication details on the runtime experiment
+6.
+Appendix F Replication details on the runtime experiment
 Apart from the theoretical runtime efficiency, we measured the actual end-to-end runtime in seconds, to better analyze the practical runtime efficiency between PINE and RoToR.
@@
 We Except for the experiments on Llama-3.1-70B-Instruct, we only use a single A6000 GPU for all of the experiments.
-A.7 Input data examples
+Appendix G Input data examples
 To illustrate the input and output formats used in our experiments, we provide example inputs for the Lost-in-the-Middle (LitM), Knowledge Graph Question Answering (KGQA), and MMLU datasets.
@@
 Figure 11: Example input for the MMLU benchmark.
-Appendix B Why is removing index bias an important detail for invariant models to be effective?
+Appendix H Why is removing index bias an important detail for invariant models to be effective?
 The alphabetic index (A/B/C/D) introduced in Fig.
@@
 For example, even for cases where the index ordering of the input was in alphabetical order (A->B->C->D), the ordering-invariant model may interpret contexts with (C->A->B->D) at one point (e.g., when the query is D on self attention), which can cause unnatural, out-of-distribution representation, leading to decreased performance.
-Appendix C Statisticial significance before and after shuffling segments
+Appendix I Statisticial significance before and after shuffling segments
 We conducted paired two-tailed t-tests (Table 8) for both the baseline (“original”) model and our proposed method (RoToR), using the results in Table 7.
@@
 While the original model shows a statistically significant performance drop with shuffled inputs, RoToR remains unaffected, demonstrating greater robustness to segment-order perturbations.
+Appendix J Application to Long-Context Inputs
+Benchmark and protocol.
+To test whether RoToR consistantly performs well to inputs with longer contexts, we extend our evaluation to the LongBench Bai et al.
+(2024)2WikiMultihopQA task, whose multi-document questions yield input lengths from 5 k to 15 k tokens.
+Because our current RoToR and PINE implementations do not yet support advanced parallelization (e.g., FlashAttention or SDPA), GPU memory becomes prohibitive beyond $\sim$10 k tokens, especially for PINE, whose memory footprint is larger than ours.
+We therefore truncate the context at 10 k tokens, which already exceeds the lengths used in our main experiments.
+Following the official LongBench guidelines [^12] Maximum generation length 32 and qa_f1_score (LongBench-E) as the evaluation metric., we segment each long context into 512-token “listwise” inputs, enabling a direct comparison with our listwise reranking pipeline.
+To ensure robustness, we test three input-order perturbations:
+  1. 1.
+    Initial: original chunk order $[1,2,3,4,5]$;
+  2. 2.
+    Reversed: $[5,4,3,2,1]$;
+  3. 3.
+    Center-flipped: first and last halves swapped, [3,2,1,5,4].
+Results.
+Table 9 reports F1 scores for the Original model variant, RoToR, and PINE with Llama 3.1-8B-Instruct and Qwen 1.5-7B-Chat.
+Across every ordering and token-length band, RoToR retains a clear advantage, while PINE fails to run when the longest (8 k+) chunks are kept.
+These findings confirm that RoToR scales to substantially longer inputs across different input-order perturbations.
+Llama 3.1-8B-Instruct Qwen 1.5-7B-Chat
+Order Method 0–4k 4–8k 8k+ Total 0–4k 4–8k 8k+ Total
+Count 25 131 144 300 23 121 156 300
+Initial (e.g., 1,2,3,4,5) Orig. 48.3 56.8 34.0 45.1 65.6 47.9 26.7 38.2
+PINE 51.0 47.6 70.2 45.1
+RoToR 59.0 52.7 41.8 48.0 75.7 47.8 31.0 41.2
+Reversed (e.g., 5,4,3,2,1) Orig. 57.0 51.5 39.0 46.0 53.4 43.3 34.2 39.3
+PINE 43.0 49.8 64.1 48.9
+RoToR 59.0 52.0 41.0 47.3 72.8 47.6 30.8 40.8
+Center flip (e.g., 3,2,1,5,4) Orig. 47.0 47.7 35.6 41.8 61.0 40.6 32.7 38.1
+PINE 46.3 49.2 70.2 43.5
+RoToR 59.0 52.5 41.5 47.8 77.1 47.3 30.9 41.0
+Table 9: F1 scores (%) on LongBench–2WikiMultihopQA with $\sim$10k-token contexts. “Count” is the number of examples per length bucket; “–” denotes out-of-memory.
+Take-aways.
+Even without specialised long-context kernels, RoToR consistently outperforms both the original reranker and PINE, and remains robust to severe input-order perturbations.
+This suggests our approach can generalize to substantially longer inputs once memory and kernel constraints are alleviated.
+Appendix K Robustness to task templates
+Llama-3.1-8B-Instruct Qwen-1.5-4B-Chat Qwen-1.5-7B-Chat
+$N=30$ $N=50$ $N=30$ $N=50$ $N=30$ $N=50$
+Method Acc. EM F1 Acc. EM F1 Acc. EM F1 Acc. EM F1 Acc. EM F1 Acc. EM F1
+Original template, Initial ordering
+Original 50.2 44.0 51.9 50.0 44.0 51.7 30.7 27.9 34.9 31.6 28.6 35.8 31.5 27.8 35.4 31.7 28.0 35.7
+PINE 51.5 45.0 52.6 51.6 45.1 52.6 31.6 28.7 35.6 31.6 28.8 35.3 32.3 28.8 36.4 32.0 28.5 35.9
+RoToR 53.1 46.5 54.1 52.9 46.0 53.6 32.0 29.0 35.7 32.7 29.6 36.2 34.3 29.8 37.7 34.3 30.1 38.0
+RoToR-MonoT5 51.6 45.0 52.5 52.4 45.4 52.8 32.3 29.1 36.2 32.3 29.3 35.9 32.9 28.4 36.3 32.9 28.9 36.6
+RoToR-Freq. 52.6 46.1 53.7 53.1 46.4 53.7 32.3 29.2 36.0 32.3 29.2 35.9 33.7 29.5 37.2 33.5 29.5 37.2
+Template-swap
+Original 50.0 44.1 51.8 50.2 44.3 51.9 31.1 27.8 34.9 31.7 28.3 35.3 31.4 27.6 35.2 32.0 28.0 35.7
+PINE 52.0 45.7 52.9 52.0 45.3 52.8 31.9 28.8 35.8 31.7 28.6 35.4 31.9 28.5 36.0 31.5 28.2 35.7
+RoToR 52.7 46.4 54.0 52.9 46.4 53.7 31.8 28.2 35.0 32.4 29.0 35.6 34.1 29.8 37.6 34.0 29.9 37.7
+RoToR-MonoT5 51.5 45.2 52.6 52.5 45.7 53.1 32.4 29.0 36.3 32.6 29.3 35.9 32.9 28.5 36.4 32.6 28.5 36.3
+RoToR-Freq. 52.3 46.2 53.7 52.9 46.5 53.8 31.9 28.4 35.4 32.4 28.8 35.6 34.0 29.9 37.4 33.8 29.6 37.4
+Table 10: Results on the Mintaka (KGQA) dataset on different models, before (top block, also reported at main paper) and after (bottom block) the template-swap.
+N refers to number of top-k segments per query.
+RoToR variants consistently outperform the Original and PINE baselines, and their performance is stable under the swapped template, indicating robustness to instruction wording.
+We evaluated whether the proposed selective-routing methods remain effective when the surrounding task instructions are re-phrased.
+Concretely, we performed a template-swap experiment on the KGQA benchmark, specifically on the initial ordering setup.
+The original prompt (Figure 10) began with
+“Below are the facts in the form of triples meaningful to answer the question.”
+and required the model to output only a JSON object. In the swapped template we replaced the first sentence with
+“Below are knowledge statements expressed as triples meaningful to answer the question.”
+leaving all other instructions unchanged.
+Table 10 reports the results.
+Across all three backbone models and both retrieval depths ($N{=}30,50$), RoToR and its variants retain similar absolute scores and continue to outperform both the Original and PINE baselines, indicating strong robustness to superficial wording changes in the task template.
+Appendix L Additional Statistics on Selective Routing Assignment
+Table 11 complements the main results in Table 3 by reporting the selection ratio, accounting for the percentage of evaluation queries for which the RoToR branch is chosen over the vanilla branch, under Selective Routing (SR).
+[^13] All figures are computed over the full evaluation set of 14,015 queries.
+We break the analysis down by (i) the global sorting strategy (Lexical, MonoT5, or Freq.), and (ii) three model backbones.
+The table distinguishes three order–based conditions:
+Llama-3.1-8B-Instr. Qwen1.5-4B-Chat Qwen1.5-7B-Chat
+Sorting Init. Rev. Avg. Init. Rev. Avg. Init. Rev. Avg.
+Lexical 7.0 8.5 $7.3\!\pm\!0.8$ 5.9 6.2 $6.2\!\pm\!0.4$ 10.3 10.6 $9.9\!\pm\!0.6$
+MonoT5 6.9 7.6 $6.7\!\pm\!1.5$ 8.0 12.5 $9.8\!\pm\!2.1$ 10.7 10.9 $10.7\!\pm\!0.7$
+Freq. 6.4 6.7 $6.9\!\pm\!0.5$ 8.5 10.9 $9.4\!\pm\!1.6$ 10.7 11.1 $11.1\!\pm\!0.8$
+Table 11: Selection ratio (%) of the RoToR variant under SR. Higher values indicate more frequent routing to RoToR.
+For table 11, Init.
+refers to the original ordering (e.g., in abcd order), Rev.
+refers to the reversed ordering (e.g., in dcba but assigned as abcd), and Avg.
+is the average selection ratio for all possible (4!-1) re-orderings, with standard deviation.
+Empirically, the RoToR model tends to be selected more frequently under reversed orderings, whereas under the original ordering, the vanilla model is chosen slightly more often.
+The exact ratio varies by model and sorting strategy.